### v1

# The local moduli of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres and invertible polynomials

Abstract.
We compute the dimension of the local moduli space of Sasaki-Einstein metrics for links of invertible polynomials coming from the list of Johnson and Kollár of anticanonically embedded Fano 3-folds of index 1 that produce $\mathbb{Q}$-homology 7-spheres, that is, 7-manifolds whose $\mathbb{Q}$-homology equals that of the 7-sphere. In order to do so, we propose additional conditions to the Diophantine equations associated to this problem. We also find solutions for the problem associated to the moduli for the Berglund-Hübsch duals of links arising from Thom-Sebastiani sums of chain and cycle polynomials. In particular we give an explicit description of local moduli of the complex structure for the corresponding Kähler-Einstein orbifold which in this setting is a ${\mathbb{Q}}$-homology projective 3-spaces with quotient singularities.
Keywords: Moduli, Sasaki-Einstein metrics, Rational homology spheres, Berglund-Hübsch.
Mathematics Subject Classification 53C25; 32G07; 32G13.
1. Introduction
The moduli problem, that is, the study of spaces that classify geometric structures on manifolds up to isomorphism, is highly relevant in Geometry. In particular the moduli of Einstein structures has received much attention, see Chapter 12 of [2]. In many cases, finding a space with suitable properties that parametrizes the isomorphism classes is too hard and even is the space is theoretically described, sometimes in practice it is not possible to provide effective tools to compute the moduli space on explicit examples. However, the local problem sometimes is easier to handle. This was the approach taken by Kodaira and Spencer in their canonical work on deformations of complex structures on compact manifolds [25].
For Sasakian manifolds, which are roughly speaking the odd dimensional analogue of Kähler manifolds, the moduli problem has been addressed and an appropriate notion of moduli space has been achieved, see [3] and references therein (c.f.[31]). In particular, finding the number of components of this moduli has been studied and used to obtain lower bounds for the dimension of the moduli space for links of Brieskorn-Pham polynomials and Smale manifolds [10, 20, 7, 26]. For instance, the moduli space of Sasakian metrics with positive Ricci curvature on odd homotopy spheres bounding parallelizable manifolds has infinitely many positive components [8]. An important ingredient to obtain information on the moduli is given by the use of invertible polynomials, particularly to describe the moduli of Sasaki-Einstein metrics.
Indeed, invertible polynomials have been studied in the context of Sasakian geometry for some time. These polynomials, thought as defining subsets of weighted projective spaces, produce links with some interesting features both at topological/differential and Riemannian level. As a matter of fact, in [4] Boyer and Galicki generalized a method introduced by Kobayashi in [22] to produce Sasaki-Einstein metrics on links of hypersurface singularities. From there, many examples of highly connected Sasaki-Einstein manifolds were found. For such examples the existence of orbifold Kähler-Einstein metrics on various Fano orbifolds is verified using different incarnations of the $\alpha$-invariant method developed by Tian, Nadel, Demailly and Kollár [35, 29, 16]. In [20, 6, 7, 11] special interest was given to 5-dimensional manifolds and homotopy spheres realized as links of Brieskorn-Pham polynomials which are an important block in the classification of invertible polynomials. In these articles the local moduli space of Sasaki-Einstein structures for these links were also studied. For instance in [7] they exhibited homotopy spheres given as links of dimension $32n-$ 3 such that the number of inequivalent families of Sasaki-Einstein metrics and the dimension of their moduli grows double exponentially with the dimension and they also showed that all the 28 oriented diffeomorphism classes on $S^{7}$ admit inequivalent families of Sasakian-Einstein structures. Later, in [12] it is shown that the moduli space of positive Sasaki classes with vanishing first Chern class for manifolds of the form $S^{2n}\times S^{2n+1}$ or for odd dimensional homotopy spheres among others, has a countably infinite number of components of dimension greater than one and these contain no extremal Sasaki metrics at all. Again this information is produced through links of Brieskorn-Pham polynomials.
In [5], based on the complete list of Johnson and Kollár of anticanonically embedded Fano 3-folds of index one [19], many $\mathbb{Q}$-homology 7-spheres admitting Sasaki-Einstein metrics were constructed using the Kobayashi-Boyer-Galicki method. In [13] the authors were able to enlarge the previous list of examples due to a result [18] on the validity of an Orlik’s conjecture [32] on how to compute the homology of links of invertible polynomials. In [14], in joint work with R. Gomez, we studied these examples in a more systematic way and showed that this list of $\mathbb{Q}$-homology 7-spheres remain $\mathbb{Q}$-homology 7-spheres under the Berglund-Hübsch transpose rule from classical mirror symmetry [1] and under very mild conditions all of them leave the Einstein condition invariant. We found that all the invertible polynomials producing $\mathbb{Q}$-homology 7-spheres and admitting Sasaki-Einstein metrics were polynomials of cycle type, or iterated Thom-Sebastiani sums of cycle and chain types and not of Brieskorn-Pham type. It is natural to determine the family of inequivalent Sasaki-Einstein on each of these $\mathbb{Q}$-homology 7-spheres. In this article we study the space of deformations of Sasaki-Einstein metrics for this sort of links. We found that only the latter case admits non-equivalent Sasaki-Einstein metrics. Actually we show that
  • For polynomials of cycle type producing $\mathbb{Q}$-homology 7-spheres the corresponding local moduli space of Sasaki-Einstein metrics has dimension zero, that is, links of this sort do not admit inequivalent families of Sasaki-Einstein structures, see Subsection 3.1.
  • For Thom-Sebastiani sums of chain and cycle polynomials the generators of local moduli space of Sasaki-Einstein metrics depend merely on a few parameters that are given in terms of the exponents of the polynomial and the decomposition of the corresponding weights, see Subsection 3.2.
  • For the Berglund-Hübsch dual link, where one only needs to study polynomials of chain-cycle type of Fano index greater than one (see Subsection 3.3), we found that $\mathbb{Q}$-homology 7-spheres given as links coming from polynomials of this type do not admit inequivalent families of Sasaki-Einstein structures with the exception of five elements.
In order to achieve these results, we find the generators of the infinitesimal deformations of the complex structures for orbifolds arising from families of polynomials of the types mentioned above (see Subsection 2.2). In particular our results give information on the connected components of the moduli of $\mathbb{Q}$-homology 7-spheres with specific differential structure which can be expressed as link of isolated hypersurface singularities coming from Johnson and Kollár’s list of Fano 3-folds of index 1. Our findings can be interpreted in two different setting:
  • Seifert $S^{1}$-bundles are $\mathbb{Q}$-homology spheres if and only if the corresponding orbifolds are $\mathbb{Q}$-homology complex projective spaces [21], so our results describes some components of the moduli space of ${\mathbb{Q}}$-homology complex projective 3-spaces with quotient singularities.
  • Sasaki-Einstein structures on the manifold determine Ricci-flat Kähler cone metrics on the corresponding affine cone [8], so our results give information on the moduli of Calabi-Yau cones [31].
The general strategy is to find the arithmetic conditions provided by the invertible polynomials $f$ that produce links $L_{f}$ that are $\mathbb{Q}$-homology 7 spheres and also admit Sasaki-Einstein metrics. Then we interpret these arithmetic conditions as constraint to obtain all possible monomials
$z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}},$
of certain degree, that describe the space of deformations of holomorphic structures on the orbifold $Z_{f}=\{f=0\}/\mathbb{C}^{*}$ associated to the link. Thus in order to find the number of monomials we obtain a Diophantine equation subject to certain the conditions that we are able to solve for the different invertible polynomials that appear in this problem.
The paper is organized as follows. Section 2 reviews the background material relevant for this paper. In Section 3 we relate the problem of finding the moduli number of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres to solving the Diophantine equations subject to certain arithmetic conditions extracted from the building blocks that conform the invertible polynomial associated to this problem; then we find the dimension of the corresponding local moduli space of Sasaki-Einstein metrics and prove Theorems 3.1.1, 3.2.1, 3.3.1.
2. Preliminaries
2.1. Sasakian structures on links and the Berglund-Hübsch rule
Below we give the basics on Sasakian geometry and list some important results that we will need, the canonical references here are [8] and [34]. Then we focus on links of invertible polynomials with certain arithmetic conditions which produce $\mathbb{Q}$-homology 7-spheres and explain how the Berglund-Hübsch transpose rule can be used in this framework.
Sasakian Geometry: Sasakian geometry is a special type of contact metric structure on a $(2n+1)$-dimensional manifold $M$ described by the quadruple of tensors $(\Phi,\xi,\eta,g)$ such that $\eta$ is a contact 1-form, $\Phi$ is an endomorphism of the tangent bundle, $g=d\eta\circ(\mathbb{I}\times\Phi)+\eta\otimes\eta$ is a Riemannian metric and $\xi$ is the Reeb vector field which is Killing. Moreover, the underlying CR-structure $\left(\mathcal{D},\left.\Phi\right|_{\mathcal{D}}\right)$ is integrable, where $\mathcal{D}=\ker\eta.$
A Sasaki manifold has a transverse Kählerian structure $(\nu(\mathcal{F}_{\xi}),\bar{J})$ determined by the normal bundle $\nu(\mathcal{F}_{\xi})$ of the characteristic foliation $\mathcal{F}_{\xi}$ determined by $\xi$ and the natural complex structure $\bar{J}$ in $\nu(\mathcal{F}_{\xi})$ induced by $\left.\Phi\right|_{\mathcal{D}}.$ In fact, when all the orbits of $\xi$ are closed, the Reeb vector field $\xi$ generates a locally free circle action whose quotient is a Kähler orbifold. In this case the Sasakian structure is called quasi-regular. When the action is free it is called regular and the quotient is a Kähler manifold. In the irregular case when the orbits of the Reeb vector field are not closed the local quotients are Kähler.
On the other hand, if we consider the cone $C(M)=M\times\mathbb{R}^{+}$ with symplectic form $\omega=$ $d\left(r^{2}\eta\right)$ where $r$ is the radial coordinate. Using the Liouville vector field $\Psi=r\partial_{r}$ we define a natural complex structure $I$ on $C(M)$ by
$IX=\Phi X+\eta(X)\Psi,\quad I\Psi=-\xi,$
where $X$ is a vector field on $M$, and $\xi$ is understood to be lifted to $C(M)$. By adding the apex of the cone we obtain an affine variety $Y=C(M)\cup\{0\}$ which has been intensely studied to obtain K-stability theorems in the manner of the work of Chen-Donaldson-Sun in the context of Sasakian manifolds [15]. We know that $(M,g)$ is Sasakian if the cone $(C(M),\omega,I,\bar{g})$ is Kähler with Kähler metric given by $\bar{g}=dr^{2}+r^{2}g.$ Moreover, the Sasakian structure corresponds to a polarized affine variety $(Y,I,\xi)$ polarized by the Reeb vector field $\xi$ and this affine variety admits a Ricci-flat Kähler cone metric if and only if $(M,g)$ admits a Sasaki-Einstein metric.
Let us denote by $\mathcal{S}(M)$ the space of all Sasakian structures on $M$ and give it the $C^{\infty}$ Fréchet topology as sections of vector bundles, and denote by $\mathcal{S}(M,\xi,\bar{J})$ the subset of Sasakian structures with Reeb vector field $R$ and transverse holomorphic structure $\bar{J}$. The set $\mathcal{S}(M,\xi,\bar{J})$ is endowed with the subspace topology. Consider $\mathcal{SE}(M)$ the subspace of Sasaki-Einstein metrics in $\mathcal{S}(M,\xi,\bar{J})$ and let $\mathfrak{Aut}(\bar{J})$ be the group of complex automorphisms of $(C(M),I)$ that commute with $\Psi-i\xi$. This group descends to an action on $(M,\mathcal{S})$ commuting with $\xi.$ Denote by $\mathfrak{Aut}(\bar{J})_{0}$ the connected component of $\mathfrak{Aut}(\bar{J}).$ Then one defines the local moduli space of Sasaki-Einstein metrics [3] by
$\mathfrak{M}^{SE}(M)=\mathcal{S}\mathcal{E}(M)/\mathfrak{Aut}(\bar{J})_{0}.$
As we will see in Subsection 2.2, this space has a very concrete description for links of invertible polynomials.
An important source of examples of Sasakian manifolds in a very general way can be given as total spaces of $S^{1}$-orbibundles over projective orbifolds. As a special case, one can build up Sasakian structures on links of hypersurface singularities of weighted homogeneous polynomials and we explain how to do this below.
Let us recall that a polynomial $f\in\mathbb{C}\left[z_{0},\ldots,z_{n}\right]$ is said to be a weighted homogeneous polynomial of degree $d$ and weight $\mathbf{w}=$ $\left(w_{0},\ldots,w_{n}\right)$, $w_{i}\in\mathbb{Z}^{+}$, and if for any $\lambda\in\mathbb{C}^{*}$
$f\left(\lambda^{w_{0}}z_{0},\ldots,\lambda^{w_{n}}z_{n}\right)=\lambda^{d}f\left(z_{0},\ldots,z_{n}\right).$
We assume that $f$ is chosen so that the affine algebraic variety
$V_{f}=\{f=0\}\subset\mathbb{C}^{n+1}$
is smooth everywhere except at the origin in $\mathbb{C}^{n+1}$ which is equivalent to assume that the weighted variety
$(V_{f}-\{0\})/\mathbb{C}^{*}\subset\mathbb{P}(\mathbf{w})$
is quasi-smooth of dimension $n.$
A link $L_{f}(\mathbf{w},d)$ of a hypersurface singularity is defined as $f^{-1}(0)\cap S^{2n+1}$, where $S^{2n+1}$ is the $(2n+1)$-sphere in Euclidean space. By the Milnor fibration theorem [27], $L_{f}(\mathbf{w},d)$ is a closed ($n-2)$-connected manifold that bounds a parallelizable manifold with the homotopy type of a bouquet of $n$-spheres. Furthermore, $L_{f}(\mathbf{w},d)$ admits a quasi-regular Sasaki structure $\mathcal{S}=\left(\xi_{\mathbf{w}},\eta_{\mathbf{w}},\Phi_{\mathbf{w}},g_{\mathbf{w}}\right)$ which is the restriction of the weighted Sasakian structure on the sphere $S^{2n+1}$ with Reeb vector field $\xi_{\mathrm{w}}=\sum_{k=0}^{n}w_{k}\left(y_{k}\partial_{{x}_{k}}-x_{k}\partial_{{y}_{k}}\right).$ Moreover, if one considers the locally free $S^{1}$-action induced by the weighted $\mathbb{C}^{*}$ action on $f^{-1}(0)$ the quotient space of the link $L_{f}(\mathbf{w},d)$ by this action is the weighted hypersurface $Z_{f},$ a Kähler orbifold: we have the following commutative diagram [8]
[…]
where $S_{\mathbf{w}}^{2n+1}$ denotes the unit sphere with a weighted Sasakian structure, $\mathbb{P}(\mathbf{w})$ is a weighted projective space coming from the quotient of $S_{\mathbf{w}}^{2n+1}$ by a weighted circle action generated from the weighted Sasakian structure. The top horizontal arrow is a Sasakian embedding and the bottom arrow is a Kählerian embedding and the vertical arrows are orbifold Riemannian submersions.
It follows from the orbifold adjunction formula that the link $L_{f}$ admits a positive Ricci curvature if the quotient orbifold $Z_{f}$ by the natural action $S^{1}$ is Fano, which is equivalent to
$I=|{\bf w}|-d_{f}>0.$
Here $|\mathbf{w}|=\sum_{i=0}^{m}w_{i}$ denotes the norm of the weight vector $\mathbf{w}$ and $d_{f}$ is the degree of the polynomial $f$, one refers to its difference $I$ as the index.
In [22], Kobayashi showed that the link of a cone over a smooth projective variety $Z\subset\mathbb{P}^{n}$ carries a natural Einstein metric if and only if $Z$ is Fano and $Z$ carries a Kähler-Einstein metric. In [4], the authors generalized this result to weighted cones and furthermore gave an algorithm, the Kobayashi-Boyer-Galicki method, to obtain $(n-1)$-connected Sasaki-Einstein $(2n+1)$-manifolds from the existence of orbifold Fano Kähler-Einstein hypersurfaces $Z_{f}$ in weighted projective $2n$-space $\mathbb{P}(\mathbf{w}).$
In order to obtain topological information of the link, an important device is given by the Alexander polynomial $\Delta_{f}(t)$ in [27] associated to a link $L_{f}$ of dimension $12n-$ 1 is the characteristic polynomial of the monodromy map $h_{*}:H_{n}(F,\mathbb{Z})\rightarrow H_{n}(F,\mathbb{Z})$ induced by the circle action on the Milnor fibre $F$. Then $\Delta_{f}(t)=\operatorname{det}\left(t{\mathbb{I}}-h_{*}\right)$. Now both $F$ and its closure $\bar{F}$ are homotopy equivalent to a bouquet of $n$-spheres $S^{n}\vee\cdots\vee S^{n},$ and the boundary of $\bar{F}$ is the link $L_{f}$, which is $(n-2)$-connected. The Betti numbers $b_{n-1}\left(L_{f}\right)=b_{n}\left(L_{f}\right)$ equal the number of factors of $(t-1)$ in $\Delta_{f}(t)$. From the Wang sequence of the Milnor fibration (see [32])
$\quad 0\longrightarrow H_{n}\left(L_{f},\mathbb{Z}\right)\longrightarrow H_{n}(F,\mathbb{Z})\stackrel{{\scriptstyle\mathbb{I}-h_{*}}}{{\longrightarrow}}H_{n}(F,\mathbb{Z})\longrightarrow H_{n-1}\left(L_{f},\mathbb{Z}\right)\longrightarrow 0$
one obtain that $L_{f}$ is a $\mathbb{Q}$-homology sphere if and only if $\Delta_{f}(1)\neq 0$ and the order of $H_{n-1}\left(L_{f},\mathbb{Z}\right)$ equals $|\Delta_{f}(1)|.$ In the case that $f$ is a weighted homogeneous polynomial there is an algorithm due to Milnor and Orlik [28] to calculate the free part of $H_{n-1}\left(L_{f},\mathbb{Z}\right).$ The authors associate to any monic polynomial $f$ with roots $\alpha_{1},\ldots,\alpha_{k}\in\mathbb{C}^{*}$ its divisor
$\operatorname{div}f=\left\langle\alpha_{1}\right\rangle+\cdots+\left\langle\alpha_{k}\right\rangle$
as an element of the integral ring $\mathbb{Z}\left[\mathbb{C}^{*}\right].$ Let $\Lambda_{n}=\operatorname{div}\left(t^{n}-1\right)$. Then the divisor of $\Delta_{f}(t)$ is given by
$\operatorname{div}\Delta_{f}=\prod_{i=0}^{n}\left(\frac{\Lambda_{u_{i}}}{v_{i}}-\Lambda_{1}\right),$
where the $u_{i}^{\prime}s$ and $v_{i}^{\prime}s$ are given terms of the degree $d$ of $f$ and the weight vector ${\bf w}=(w_{0},\ldots w_{n})$ by the equations
$u_{i}=\frac{d}{\operatorname{gcd}\left(d,w_{i}\right)},\quad v_{i}=\frac{w_{i}}{\operatorname{gcd}\left(d,w_{i}\right)}.$
Invertible polynomials and the Berglund-Hübsch transpose: In [13, 14] we used the Kobayashi-Boyer-Galicki method to establish the existence of Sasaki-Einstein metrics on links of invertible polynomials. Recall that a polynomial $f\in\mathbb{C}\left[x_{1},\ldots,x_{n}\right]$ is invertible if it is of the form
$f=\sum_{i=1}^{n}\prod_{j=1}^{n}x_{j}^{a_{ij}}$
where $A=\left(a_{ij}\right)_{i,j=1}^{n}$ is a non-negative integer-valued matrix satisfying:
  1. (1)
    $A$ is invertible over $\mathbb{Q}$;
  2. (2)
    $f$ is quasi-homogeneous, i.e., there exists positive integers $w_{j}$ such that $d:=\sum_{j=1}^{n}w_{j}a_{ij}$ is constant for all $i$; and
  3. (3)
    $f$ is quasi-smooth, i.e., $f:\mathbb{C}\left[x_{1},\ldots,x_{n}\right]\rightarrow\mathbb{C}$ has exactly one critical point at the origin.
Due to the Kreuzer-Skarke classification of invertible polynomials [25] we know that any invertible polynomial, up to permutation of variables, can be written as a Thom-Sebastiani sum of three types of polynomials usually called atoms:
  1. (1)
    Fermat type: $w=x^{r}$,
  2. (2)
    Chain type: $w=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}x_{3}+\ldots+x_{n-1}^{a_{n-1}}x_{n}+x_{n}^{a_{n}}$, and
  3. (3)
    Loop or cycle type: $w=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}x_{3}+\ldots+x_{n-1}^{a_{n-1}}x_{n}+x_{n}^{a_{n}}x_{1}$.
For instance Brieskorn-Pham polynomials $f=x_{1}^{a_{1}}+x_{2}^{a_{2}}+\ldots+x_{n}^{a_{n}}$ are Thom-Sebastiani sums of Fermat type polynomials.
In [5], [13] all the invertible polynomials that produce $\mathbb{Q}$-homology 7-spheres and admit Sasaki-Einstein metrics are polynomials of cycle type, chain type, or iterated Thom-Sebastiani sums of these types. These polynomials were taken from the list of Johnson and Kollár of anticanonically embedded Fano 3-folds of index 1 [19]. All these polynomials satisfy one of the following conditions:
  • The weights and the degree satisfy $\operatorname{gcd}(d,w_{i})=1$ for all $i=0,\dots 4,$ which leads to singularities of cycle type.
  • The weights are subject to $(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{2}v_{0},m_{2}v_{1},m_{2}v_{2},m_{3}v_{3},m_{3}v_{4})$ with $\gcd(m_{2},m_{3})=1$ and $m_{2}m_{3}=d,$ which leads to singularities that can described as iterated Thom-Sebastiani sums of chain, cycle type and Fermat singularities. More precisely, the types of singularities obtained have the following form:
    Type I (Fermat-Cycle): $z_{0}^{a_{0}}+z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
    Type II (Chain-Cycle): $z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
    Type III (Cycle-Cycle): $z_{1}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}.$
Proofs of the following statements can be found in [5] and/or in the proof of Theorem 3.2 in [13]:
  1. (1)
    Consider the link $L(\textbf{w},d)$ of a weighted homogeneous polynomial $f$ with $\operatorname{gcd}(d,w_{i})=1$ for all $i=0,\dots 4,$ then the Milnor number $m(L_{f})$ for $L_{f}$ is given by
    $m(L_{f})+1=d(b_{n-1}+1)\,\,\mathrm{and}\,\,H_{3}\left(L_{f},\mathbb{Z}\right)_{tor}=\mathbb{Z}_{d}.$
    In particular, if $f$ is given as a polynomial of cycle type
    $f=z_{4}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{1}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
    of degree $d$ in the projective space $\mathbb{P}(\textbf{w})$, where $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ then from the equations
    $a_{0}w_{0}+w_{4}=d,\quad a_{1}w_{1}+w_{0}=d,$
    $\quad a_{2}w_{2}+w_{1}=d,\quad a_{3}w_{3}+w_{2}=d,\quad a_{4}w_{4}+w_{3}=d$
    we obtain
    $a_{0}a_{1}a_{2}a_{3}a_{4}=\left(\frac{d-w_{4}}{w_{0}}\right)\left(\frac{d-w_{0}}{w_{1}}\right)\left(\frac{d-w_{1}}{w_{2}}\right)\left(\frac{d-w_{2}}{w_{3}}\right)\left(\frac{d-w_{3}}{w_{4}}\right)=m\left(L_{f}\right).$
    So in the case the link is a $\mathbb{Q}$-homology 7-sphere we obtain
    $a_{0}a_{1}a_{2}a_{3}a_{4}=d-1.$ (2.1.1)
  2. (2)
    If the link $L(\textbf{w},d)$ is such that its weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfies $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{2}v_{0},m_{2}v_{1},m_{2}v_{2},m_{3}v_{3},m_{3}v_{4})$, $\gcd(m_{2},m_{3})=1$ and $m_{2}m_{3}=d.$ One obtains the equality
    $\operatorname{div}\Delta_{f}=\alpha({\bf w})\beta({\bf w})\Lambda_{d}+\beta({\bf w})\Lambda_{m_{3}}-\alpha({\bf w})\Lambda_{m_{2}}-\Lambda_{1},$
    with the two positive integers $\alpha(\textbf{w})$ and $\beta(\textbf{w})$ depending on the weights:
    $\alpha(\textbf{w})=\frac{m_{2}}{v_{3}v_{4}}-\frac{1}{v_{3}}-\frac{1}{v_{4}}$ (2.1.2)
    and
    $\beta(\textbf{w})=\left(\frac{m_{3}}{v_{0}v_{1}}-\frac{1}{v_{1}}-\frac{1}{v_{0}}\right)\left(\frac{m_{3}}{v_{2}}-1\right)+\frac{1}{v_{2}}.$ (2.1.3)
    It is known that if the link is a $\mathbb{Q}$-homology sphere, then $\beta({\bf w})=1.$
    Furthermore, if $f$ has a cycle block of the form
    $z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}},$
    then
    $a_{2}w_{2}+w_{4}=d,\quad a_{3}w_{3}+w_{2}=d\text{ and }a_{4}w_{4}+w_{3}=d.$
    From the assumptions on the weight vector $\mathbf{w}$ we have
    $a_{2}a_{3}a_{4}$ $=\left(\frac{d-w_{4}}{w_{2}}\right)\left(\frac{d-w_{2}}{w_{3}}\right)\left(\frac{d-w_{3}}{w_{4}}\right)$
    $=\left(\frac{m_{3}-v_{4}}{v_{2}}\right)\left(\frac{m_{3}-v_{2}}{v_{3}}\right)\left(\frac{m_{3}-v_{3}}{v_{4}}\right)$
    $=\frac{m_{3}^{3}-\left(v_{2}+v_{3}+v_{4}\right)m_{3}^{2}+\left(v_{2}v_{3}+v_{2}v_{4}+v_{3}v_{4}\right)m_{3}-v_{2}v_{3}v_{4}}{v_{2}v_{3}v_{4}}$
    $=m_{3}\left(\frac{m_{3}^{2}-\left(v_{2}+v_{3}+v_{4}\right)m_{3}+v_{2}v_{3}+v_{2}v_{4}+v_{3}v_{4}}{v_{2}v_{3}v_{4}}\right)-1.$
    So if the corresponding link is a $\mathbb{Q}$-homology sphere, it follows that
    $\beta(\mathbf{w})=\frac{m_{3}^{2}-\left(v_{2}+v_{3}+v_{4}\right)m_{3}+v_{2}v_{3}+v_{2}v_{4}+v_{3}v_{4}}{v_{2}v_{3}v_{4}}=1.$
    Substituting this equality in the previous equation, we obtain
    $a_{2}a_{3}a_{4}+1=m_{3}.$ (2.1.4)
Recall that the Berglund-Hübsch transpose rule considers an invertible polynomial $f=\sum_{j=0}^{n}\prod_{i=0}^{n}x_{i}^{a_{ij}}$ cutting out an orbifold of degree $d$ in $\mathbb{P}(\bf w)$ and defines the transpose polynomial $f^{T}$ by transposing the exponential matrix $A=\left(a_{ij}\right)_{i,j}$ of the original polynomial, that is,
$f^{T}=\sum_{i=1}^{n}\prod_{j=1}^{n}x_{j}^{a_{ji}},$
again an invertible polynomial that cuts out an orbifold of degree $\tilde{d}$ in $\mathbb{P}(\tilde{\bf w}).$ Then one considers the links $L_{f}(\mathbf{w},d)$ and $L_{f^{T}}(\tilde{\mathbf{w}},\tilde{d})$ associated to each of these polynomials. We will sometimes say that these two links are Berglund-Hübsch duals to one another. The following diagram succinctly summarizes the procedure described above, where BH denotes the Berglund-Hübsch transpose rule:
${{f=0}}$${{f^{T}=0}}$${L_{f}(\mathbf{w},d)}$${L_{f^{T}}(\tilde{\mathbf{w}},\tilde{d}).}$$\scriptstyle{BH}$$\scriptstyle{BH}$
In [14] the Berglund-Hübsch transpose rule is used to produce links with the $\mathbb{Q}$-homology of a 7-sphere and it was found that this rule only produces twins for singularities of cycle type or of type I and type III. Recall [5, 14] that two links $L_{f}$ and $L_{g}$ of an isolated hypersurface singularity are called twins if both are $\mathbb{Q}$-homology $(2n+1)$-spheres and they satisfy $m\left(L_{f}\right)=m\left(L_{g}\right),d_{g}=d_{f}$, and $H_{n}\left(L_{f},\mathbb{Z}\right)=H_{n}\left(L_{g},\mathbb{Z}\right)$. However, for case II, the Berglund-Hübsch transpose rule does not preserve neither torsion nor Milnor number. Its dual links will receive special attention is Subsection 3.3.
2.2. Deformations of transverse holomorphic Sasakian structures
Locally the moduli space of Sasaki isotopy classes is determined by the deformation theory of the transverse holomorphic structure of the foliation $\mathcal{F}.$ So the usual thing to do is to fix the contact structure and deform the transverse holomorphic structure via Kodaira-Spencer theory.
The cohomological groups with respect to the sheaf of holomorphic vector fields $T_{X}$ (on a complex manifold $X$) have precise interpretation in terms of deformation theory [23]: $H^{1}\left(X,T_{X}\right)$ can be seen as the space of infinitesimal deformations of the complex structure of $X,H^{2}\left(X,T_{X}\right)$ is the space of obstructions to lifting an infinitesimal deformation to an actual deformation of the complex structure of $X$. The group $H^{0}\left(X,T_{X}\right)$ can be considered as the Lie algebra of the group of holomorphic automorphisms of $X$, and if $H^{0}\left(X,T_{X}\right)=0$ then the Kuranishi family is universal. We have analogs to these notions for the transverse holomorphic case, we give a brief review of these notions following [17] for the case of interest: a Sasakian manifold $(M,\xi,\eta,\Phi,g)$ with transverse holomorphic foliation $\mathcal{F}_{\xi}.$
A germ of a deformation of a transverse holomorphic foliation $\mathcal{F}_{\xi}$ on $M$ with base space $(B,0)$ is given by an open cover $\left\{U_{\alpha}\right\}$ of $M$ and a family of local submersions $f_{\alpha,t}:U_{\alpha}\rightarrow\mathbb{C}^{n}$ parametrized by $(B,0)$ that are holomorphic in $t\in B$ for each $x\in U_{\alpha}$. For $\Theta_{\mathcal{F}_{\xi}}$ denoting the sheaf of transversely holomorphic vector fields on $M$, we have a Kodaira-Spencer map $\rho:T_{0}B\rightarrow H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ that sends $\frac{\partial}{\partial t}$ to certain class in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ defined by a section $\theta_{\alpha,\beta}$ of the sheaf $\Theta_{\mathcal{F}_{\xi}}\mid U_{\alpha}\cap U_{\beta}$. One can consider the full cohomology ring $H^{*}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, these were proven to be finite dimensional.
In [17] it is shown that there is a versal Kuranishi space of deformations given by the map $\Phi:U\rightarrow H^{2}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, for $U$ open set in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, here, as in the complex case, the base of parametrizations is given by $\Phi^{-1}(0)$. We have that if $H^{2}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)=0$, then the Kuranishi family of deformations of $\mathcal{F}_{\xi}$ is isomorphic to an open set in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$. Otherwise, the Kuranishi space may be singular. For a quasi-regular Sasakian structure $\mathcal{S}=(\xi,\eta,\Phi,g)$ with quotient orbifold $\mathcal{Z}$ one obtains the following sequence [8]
$0\longrightarrow H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)\longrightarrow H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)\longrightarrow H^{0}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)\longrightarrow H^{2}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$
where $\Theta_{\mathcal{Z}}$ denotes the sheaf of germs of holomorphic vector fields on the complex orbifold $\mathcal{Z}$. Thus the deformation of the transverse holomorphic structures of the foliation can be understood in terms of the deformation of the complex structure of the orbifold, which are described by $H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$ and in terms of the deformations of the Reeb vector in $H^{0}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$ which are described by the Sasaki cone [9]. Of course, under the absence of automorphism on the orbifold we obtain an isomorphisms between $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ and $H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right).$
Also, as shown in [30], deforming the transversely holomorphic structure may not stay Sasakian unless the $(0,2)$ component of the basic Euler class $\left[d\eta^{0,2}\right]\in H^{(0,2)}_{B}(M,\mathcal{F}_{\xi})$ is zero, where $H^{0,2}_{B}(M,\mathcal{F}_{\xi})$ is the basic Dolbeault cohomology for the transversal complex structure. In case the Sasakian structure is positive, which is the case we are interested in this article, it has been proven in [30] that $H^{0,q}_{B}(M,\mathcal{F}_{\xi})=0$ for $q>0.$
We pay particular attention to the local moduli of quasi-regular Sasakian structures on links of isolated singularities, so our focus is on orbifolds that are quasi-smooth weighted hypersurface in certain weighted projective space. For this kind of orbifolds the group $H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$ has a precise description. The proof of the next two theorems can be found in [8]. (See also [7].)
Theorem 2.2.1.
Let $Z_{f}$ be a quasi-smooth weighted hypersurface in $\mathbb{CP}(\mathbf{w})$ corresponding to the weighted homogenous polynomial $f$ of degree $d$ and weight vector $\mathbf{w}=\left(w_{0},\ldots,w_{n}\right)$. Let $\mathfrak{Aut}\left(Z_{f}\right)$ denote the group of automorphisms of $Z_{f}$. Assume also that $n\geqslant 3$. Then the complex orbifolds $\mathcal{Z}=\left(Z_{f},\mathcal{U}\right)$ form a continuous family of complex dimension
$h^{0}(\mathbb{CP}(\mathbf{w}),\mathcal{O}(d))-\sum_{i}h^{0}\left(\mathbb{P}(\mathbf{w}),\mathcal{O}\left(w_{i}\right)\right)+\operatorname{dim}\mathfrak{Aut}\left(Z_{f}\right).$ (2.2.1)
Furthermore, if the index $I=|\mathbf{w}|-d>0$ and $Z_{f}$ admits a Kähler-Einstein metric for a generic $f$ then it admits a $2\left[h^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(d))-\sum_{i}h^{0}\left(\mathbb{P}(\mathbf{w}),\mathcal{O}\left(w_{i}\right)\right)+\operatorname{dim}\mathfrak{Aut}\left(Z_{f}\right)\right]$ dimensional family of Kähler-Einstein metrics up to homothety.
Here $h^{0}=\dim_{\mathbb{C}}H^{0}$ denotes the complex dimension of the corresponding space of holomorphic sections and $\mathfrak{Aut}\left(Z_{f}\right)$ denotes the group of complex automorphisms of the orbifold $Z_{f}.$
For quasi-regular Sasaki- Einstein metrics we have the following
Theorem 2.2.2.
Let $M$ be a smooth compact manifold and let $\mathcal{S}=\left(\xi,\eta,\Phi_{t},g_{t}\right)$ be a family of quasi-regular Sasaki-Einstein structures on $M$ induced by a continuous family of inequivalent complex orbifolds $\mathcal{Z}_{t}$ with Kähler-Einstein metrics. Then the metrics $g_{t}$ are inequivalent as Einstein metrics.
From these two theorems one can compute the dimension of the local moduli of Sasaki-Einstein metrics for certain links and therefore obtain lower bounds to the moduli space of Einstein metrics on many manifolds. We will do this for $\mathbb{Q}$-homology 7-spheres that are obtained as links of invertible polynomials. The key is to consider all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $d$ which is equivalent to find all the solutions of the following Diophantine equation
$w_{0}x_{0}+w_{1}x_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d$
with variables $x_{i}\in\mathbb{Z}_{0}^{+}$ and where at least one of them is nonzero and subject to certain arithmetic constraints given by the type of invertible polynomials under study.
3. Local moduli of Sasaki-Einstein metrics on $\mathbb{Q}$-homology 7-spheres
In this section we will compute the dimension of the local moduli of Sasaki-Einstein metrics of links that are $\mathbb{Q}$-homology 7-spheres. For that we will study the space of deformations of polynomials of cycle type, chain type, or certain iterated Thom-Sebastiani sums of these types. We only focus on these types of singularities for the reasons that were argued previously in Section 2.1: if a link $L_{f}$ is a $\mathbb{Q}$-homology 7-sphere and admits Sasaki-Einstein metric then $f$ has to be one of the aforementioned types. Since Sasaki-Einstein structures on the manifold determine Ricci-flat Kähler cone metrics on the corresponding affine cone, our results give information on the moduli of Calabi-Yau cones, see Problem 3.22 in [31] where the notion of K-moduli of Fano cones was introduced.
3.1. $\mathbb{Q}$-homology 7-spheres: the cycle type
Let $f$ be an invertible polynomial of the form
$f=z_{4}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{1}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$ (3.1.1)
of degree $d$ and associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$, which is an element of Kollar’s list, such that $\gcd(d,w_{i})=1$ and $d=w_{0}+w_{1}+w_{2}+w_{3}+w_{4}-1$. From (3.1.1), we have the following relation between the weights $w_{i}$’s and the degree:
$a_{0}w_{0}+w_{4}=d,\ \ a_{1}w_{1}+w_{0}=d,\ \ a_{2}w_{2}+w_{1}=d,\ \ a_{3}w_{3}+w_{2}=d,\ \ a_{4}w_{4}+w_{3}=d$ (3.1.2)
Moreover, if the link $L_{f}$ is a $\mathbb{Q}$-homology sphere, it verifies that $d=1+a_{0}a_{1}a_{2}a_{3}a_{4}$. Using this, we can express each weight $w_{i}$ as:
$w_{i}=1-a_{i+1}+a_{i+1}a_{i+2}-a_{i+1}a_{i+2}a_{i+3}+a_{i+1}a_{i+2}a_{i+3}a_{i+4}$ (3.1.3)
where the subscript is taking mod $5$. Furthermore, considering the relations given in (3.1.2) and the fact that $\gcd(d,w_{i})=1$, we conclude that two consecutive weights $w_{i}$ and $w_{i+1}$ are always co-primes.
Now, we find the generators of the space of deformations of the orbifold $Z_{f}$. In [7], it was proven that the automorphism group of any orbifold $Z_{f}$ is finite as long as $w_{i}<\frac{1}{2}d$ for all but one of the $w_{i}^{\prime}s$ for $f$ quasi-smooth. Since the index $I=|\textbf{w}|-d=1$, the weight vector $\bf w$ does not admit polynomials that contain blocks of the form $z_{i}^{2}+z_{j}^{2}$. Thus for our $f$, $\dim\mathfrak{Aut}(Z_{f})=0,$ see [7]. Therefore, the complex dimension of $Z_{f}$ is given by
$\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}({d}))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i})).$ (3.1.4)
Let us compute $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}({d}))$. Here we consider all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $d$. Notice that it is equivalent to solve the following Diophantine equation
$w_{0}x_{0}+w_{1}x_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d$ (3.1.5)
with variables $x_{i}\in\mathbb{Z}_{0}^{+}$ and where at least one of them is nonzero.
First recall well-known result for Diophantine equations that will be used in the arguments that follow
Lemma 3.1.1.
We consider the Diophantine equation in variables $x,y$:
$ax+by=c$
If $\gcd(a,b)=1$ and $(x_{0},y_{0})$ is a particular solution, then all the solutions of the Diophantine equation are given by
$(x,y)=(x_{0}-bk,y_{0}+ak),\ \ \mbox{ where }k\in\mathbb{Z}.$
The next lemma determines all solutions of Equation (3.1.5).
Lemma 3.1.2.
Let $\bf w$ defined as in (3.1.2). Then the Diophantine equation (3.1.5) has exactly five solutions. These solutions are
$(a_{0},0,0,0,1),(1,a_{1},0,0,0),(0,1,a_{2},0,0),(0,0,1,a_{3},0)\mbox{ and }(0,0,0,1,a_{4}).$
Thus, the set of generators of the space $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by
$\{z_{0}^{a_{0}}z_{4},z_{0}z_{1}^{a_{1}},z_{1}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}}\}.$
Proof.
Since the polynomial $f$ associated to $\bf w$ is cycle, we can assume, without loss of generality, that $w_{4}=\min_{i}w_{i}$.
Now, in Equation (3.1.5), as $\gcd(w_{3},w_{4})=1$, we can define a new variable $y_{3}=w_{3}x_{3}+w_{4}x_{4}\geqslant 0$. Replacing this in Equation (3.1.5), we obtain a new Diophantine equation
$w_{0}x_{0}+w_{1}x_{1}+w_{2}x_{2}+y_{3}=d.$
Here, we define the variable $y_{2}=w_{2}x_{2}+y_{3}\geqslant 0$. Putting this in the equation above, we get
$w_{0}x_{0}+w_{1}x_{1}+y_{2}=d.$
Repeating the previous process, we define the new variable $y_{1}=w_{1}x_{1}+y_{2}\geqslant 0$. Again, replacing this above, we obtain
$w_{0}x_{0}+y_{1}=d.$ (3.1.6)
From (3.1.2), we obtain a particular solution $(a_{0},w_{4})$ of Equation (3.1.6). Thus, the general solution is given by
$x_{0}=a_{0}-t\ \ \mbox{ and }\ \ y_{1}=w_{4}+tw_{0},\ \ \mbox{ where }t\in\mathbb{Z}.$
Now, we solve the Diophantine equation
$w_{1}x_{1}+y_{2}=y_{1}.$ (3.1.7)
Since $y_{1}=w_{4}+tw_{0}$ and $w_{0}=d-a_{1}w_{1}$, we obtain
$w_{1}x_{1}+y_{2}=y_{1}=w_{4}+t(d-a_{1}w_{1})=-ta_{1}w_{1}+w_{4}+td.$
Then we have a particular solution $(-ta_{1},w_{4}+td)$. Therefore, the general solution of (3.1.7) is given by
$x_{1}=-ta_{1}-s\ \ \mbox{ and }\ \ y_{2}=w_{4}+td+sw_{1},\ \ \mbox{where }t,s\in\mathbb{Z}$
On the other hand, to solve the Diophantine equation
$w_{2}x_{2}+y_{3}=y_{2}$ (3.1.8)
we take into account that $y_{2}=w_{4}+td+sw_{1}$ and $w_{1}=d-a_{2}w_{2}$. Thus, we obtain
$w_{2}x_{2}+y_{3}=y_{2}=w_{4}+td+sw_{1}=-sa_{2}w_{2}+w_{4}+(t+s)d$
Then, a particular solution of equation above is $(-sa_{2},w_{4}+(t+s)d).$ So the general equation is
$x_{2}=-sa_{2}-r\ \ \mbox{ and }\ \ y_{3}=w_{4}+(t+s)d+rw_{2}\ \ \mbox{ where }r,s,t\in\mathbb{Z}$
Finally, we will solve the Diophantine equation
$w_{3}x_{3}+w_{4}x_{4}=y_{3}$ (3.1.9)
Since $y_{3}=w_{4}+(t+s)d+rw_{2}$ and $w_{2}=d-a_{3}w_{3}$, we have
$w_{3}x_{3}+w_{4}x_{4}=y_{3}=w_{4}+(t+s)d+rw_{2}=(t+s+r-ra_{3})w_{3}+(1+a_{4}(t+s+r))w_{4}$
This implies that a particular solution of (3.1.9) is $(t+s+r-ra_{3},1+a_{4}(t+s+r))$. Then, the general solution is given by
$x_{3}=t+s+r-ra_{3}-qw_{4}\ \ \mbox{ and }\ \ x_{4}=1+(t+s+r)a_{4}+qw_{3}\ \mbox{ where }q,r,s,t\in\mathbb{Z}$
Considering all of the above, we have the general solution for (3.1.5):
$(x_{0},x_{1},x_{2},x_{3},x_{4})=(a_{0}-t,-ta_{1}-s,-sa_{2}-r,t+s+r-ra_{3}-qw_{4},1+(t+s+r)a_{4}+qw_{3})$ (3.1.10)
where $q,r,s,t\in\mathbb{Z}$. Notice that to solve (3.1.5), we require $x_{i}\geqslant 0$ for all $i$. This implies that
$t\leqslant a_{0},\ \ \ s\leqslant-ta_{1},\ \ \ r\leqslant-sa_{2},\ \ \ ra_{3}+qw_{4}\leqslant t+s+r\ \ \mbox{ and }\ \ -qw_{3}-1\leqslant(t+s+r)a_{4}.$ (3.1.11)
On the other hand, for the variables $y_{i}$’s, we have
$(y_{1},y_{2},y_{3})=(w_{4}+tw_{0},w_{4}+td+sw_{1},w_{4}+td+sd+rw_{2}),\ \ \mbox{ where }r,s,t\in\mathbb{Z}$ (3.1.12)
As $x_{i}\geqslant 0$, then we have $y_{i}\geqslant 0$. So we obtain the following inequalities
$tw_{0}+w_{4}\geqslant 0,\ \ \ w_{4}+td+sw_{1}\geqslant 0\ \ \mbox{ and }\ \ w_{4}+td+sd+rw_{2}\geqslant 0.$ (3.1.13)
Claim 1: $q\in\{-1,0\}$. First, we prove that $q\leqslant 0$. Indeed, from (3.1.11) we have
$x_{3}=t+s+r-ra_{3}-qw_{4}\geqslant 0$ $\Rightarrow t(1-a_{3})+s(1-a_{3})+r(1-a_{3})\geqslant qw_{4}-ta_{3}-sa_{3}$
$\Rightarrow t+s+r\leqslant\dfrac{ta_{3}+sa_{3}-qw_{4}}{a_{3}-1}.$
Also, from (3.1.11), we obtain
$t+s+r\geqslant\dfrac{-qw_{3}-1}{a_{4}}.$
From the two inequalities above, we get
$\dfrac{-qw_{3}-1}{a_{4}}\leqslant\dfrac{ta_{3}+sa_{3}-qw_{4}}{a_{3}-1}\Rightarrow-qa_{3}w_{3}-a_{3}+qw_{3}+1\leqslant ta_{3}a_{4}+sa_{3}a_{4}-qa_{4}w_{4}.$
Using (3.1.2), we have $a_{3}w_{3}=d-w_{2}$ and $a_{4}w_{4}=d-w_{3}$. Replacing these in the previous inequality and simplifying, we obtain
$qw_{2}\leqslant a_{3}-1+ta_{3}a_{4}+sa_{3}a_{4}.$ (3.1.14)
Considering the equality (3.1.3) for $w_{2}$, we have
$qw_{2}$ $\leqslant a_{3}-1+a_{0}a_{3}a_{4}-ta_{1}a_{3}a_{4}$
$=-1+a_{3}-a_{3}a_{4}+a_{0}a_{3}a_{4}-a_{0}a_{1}a_{3}a_{4}+(a_{0}-t)a_{1}a_{3}a_{4}+a_{3}a_{4}$
$=-w_{2}+(a_{0}-t)a_{1}a_{3}a_{4}+a_{3}a_{4}.$
On the other hand, since $\gcd(w_{0},w_{4})=1$ and $w_{4}$ is the minimum of all $w_{i}$’s, we have $w_{4}<w_{0}$. Moreover, using (3.1.13), we obtain $t\geqslant-\dfrac{w_{4}}{w_{0}}>-1$. Thus, we have $t\geqslant 0$. Thus
$qw_{2}\leqslant-w_{2}+(a_{0}-t)a_{1}a_{3}a_{4}+a_{3}a_{4}\leqslant-w_{2}+a_{0}a_{1}a_{3}a_{4}+a_{3}a_{4}$
which implies
$q\leqslant-1+\dfrac{a_{0}a_{1}a_{3}a_{4}+a_{3}a_{4}}{w_{2}}.$ (3.1.15)
Now, we will show that $2w_{2}>a_{0}a_{1}a_{3}a_{4}+a_{3}a_{4}$. Indeed, we have
$2w_{2}>a_{0}a_{1}a_{3}a_{4}+a_{3}a_{4}$ $\Longleftrightarrow 2(1-a_{3}+a_{3}a_{4}-a_{0}a_{3}a_{4}+a_{0}a_{1}a_{3}a_{4})>a_{0}a_{1}a_{3}a_{4}+a_{3}a_{4}$
$\Longleftrightarrow a_{0}a_{3}a_{4}(a_{1}-2)+a_{3}(a_{4}-2)+2>0.$
From Inequality (3.1.15) we have
$q\leqslant-1+\dfrac{a_{0}a_{1}a_{3}a_{4}+a_{3}a_{4}}{w_{2}}<-1+2=1.$
Therefore, we have $q\leqslant 0$.
Next, we will show that $q\geqslant-1$. First, we will prove that $d<2a_{1}w_{1}$. We remember that $d=1+a_{0}a_{1}a_{2}a_{3}a_{4}$ and using (3.1.3) we write $w_{1}=1-a_{2}+a_{2}a_{3}-a_{2}a_{3}a_{4}+a_{0}a_{2}a_{3}a_{4}$. Then
$d<2a_{1}w_{1}$ $\Longleftrightarrow 1+a_{0}a_{1}a_{2}a_{3}a_{4}<2a_{1}(1-a_{2}+a_{2}a_{3}-a_{2}a_{3}a_{4}+a_{0}a_{2}a_{3}a_{4})$
$\Longleftrightarrow 1<2a_{1}-2a_{1}a_{2}+2a_{1}a_{2}a_{3}-2a_{1}a_{2}a_{3}a_{4}+a_{0}a_{1}a_{2}a_{3}a_{4}$
$\Longleftrightarrow 1<2a_{1}+2a_{1}a_{2}(a_{3}-1)+a_{1}a_{2}a_{3}a_{4}(a_{0}-2).$
Since $a_{3}\geqslant 2$ and $a_{0}\geqslant 2$, we conclude that $d<2a_{1}w_{1}$. Now, from (3.1.11) and (3.1.13), we have
$sa_{2}w_{2}\leqslant-rw_{2}\leqslant w_{4}+td+sd.$
Replacing $a_{2}w_{2}=d-w_{1}$ in the above inequality, we obtain $-sw_{1}\leqslant w_{4}+td$. Moreover, as $d<2a_{1}w_{1}$ and $w_{4}=\min_{i}w_{i}\leqslant w_{1}$, we get
$-sw_{1}\leqslant w_{1}+td<w_{1}+2ta_{1}w_{1}\Rightarrow-s<1+2ta_{1}\Rightarrow-s\leqslant 2ta_{1}.$
By (3.1.11), $a_{0}\geqslant t$, then we have
$-s\leqslant 2ta_{1}\leqslant 2a_{0}a_{1}.$ (3.1.16)
Also, from (3.1.11) we have $1+(t+s+r)a_{4}+qw_{3}\geqslant 0$, $-sa_{2}\geqslant r$ and $t\leqslant a_{0}$. Then
$-qw_{3}\leqslant 1+a_{4}(t+s+r)\leqslant 1+a_{4}(t-s(a_{2}-1))\leqslant 1+a_{4}(a_{0}-s(a_{2}-1)).$
Replacing (3.1.16) in the above inequality, we obtain
$-qw_{3}\leqslant 1+a_{4}(a_{0}+2a_{0}a_{1}(a_{2}-1))=1+a_{0}a_{4}-2a_{0}a_{1}a_{4}+2a_{0}a_{1}a_{2}a_{4}.$
Adding $01+a_{0}a_{4}-2a_{4}=1+a_{4}(a_{0}-2)>$ 0 to the right of the last inequality, we have
$-qw_{3}<2(1-a_{4}+a_{4}a_{0}-a_{4}a_{0}a_{1}+a_{4}a_{0}a_{1}a_{2})=2w_{3}\Rightarrow q>-2\Rightarrow q\geqslant-1.$
So $q\in\{-1,0\}$.
Next, we will determine the number of solutions for the two values of $q$.
  • a)
    If $q=0$, we have the solution $(a_{0},0,0,0,1).$ Indeed from (3.1.11) we have
    $t\leqslant a_{0},\ \ \ s\leqslant-ta_{1},\ \ \ r\leqslant-sa_{2},\ \ \ ra_{3}\leqslant t+s+r\ \ \mbox{ and }\ \ -1\leqslant(t+s+r)a_{4}.$ (3.1.17)
    Since $t+s+r-ra_{3}\geqslant 0$, we obtain $(t+s+r)(1-a_{3})\geqslant-sa_{3}-ta_{3}$ which implies
    $t+s+r\leqslant\dfrac{a_{3}(s+t)}{a_{3}-1}.$ (3.1.18)
    On the other hand, from the last inequality in (3.1.17), we have $t+s+r\geqslant-\dfrac{1}{a_{4}}>-1$ which means $t+s+r\geqslant 0$. So from Inequality (3.1.18) we have $s+t\geqslant 0$.
    Now from $\gcd(w_{4},w_{0})=1$ and $w_{4}=\min_{i}w_{i}$, we have $w_{4}<w_{0}$. Then $-1<-\dfrac{w_{4}}{w_{0}}\leqslant t\leqslant a_{0}$. So we obtain $0\leqslant t\leqslant a_{0}$. In addition, from (3.1.17) we have $s\leqslant-ta_{1}\leqslant 0$. As $-t\leqslant s$, we have
    $s\leqslant-ta_{1}\leqslant sa_{1}\Rightarrow 0\leqslant s(a_{1}-1)\Rightarrow s\geqslant 0.$
    Since $s\leqslant 0,$ we obtain $s=0$ so $t=0$. Finally, from (3.1.17) we have $r\leqslant-sa_{2}=0$ and since $t=s=0$, we have $r=t+s+r\geqslant 0$. Thus, $r=0$. Then, we have a solution for (3.1.5): $(a_{0},0,0,0,1)$.
  • b)
    If $q=-1$, we have to analyze for cases: In (3.1.11) we have
    $t\leqslant a_{0},\ \ \ s\leqslant-ta_{1},\ \ \ r\leqslant-sa_{2},\ \ \ ra_{3}-w_{4}\leqslant t+s+r\ \ \mbox{ and }\ \ w_{3}-1\leqslant(t+s+r)a_{4}.$ (3.1.19)
    Claim 2: $t\in\{a_{0}-1,a_{0}\}$. Indeed, from (3.1.3) we write $w_{3}=1-a_{4}+a_{4}a_{0}-a_{4}a_{0}a_{1}+a_{4}a_{0}a_{1}a_{2}$. Replacing this in the last inequality in (3.1.19) we obtain
    $t+s+r\geqslant\dfrac{w_{3}-1}{a_{4}}=\dfrac{-a_{4}+a_{4}a_{0}-a_{4}a_{0}a_{1}+a_{4}a_{0}a_{1}a_{2}}{a_{4}}=-1+a_{0}-a_{0}a_{1}+a_{0}a_{1}a_{2}.$ (3.1.20)
    From (3.1.19), using $a_{0}-t\geqslant 0$ and $-sa_{2}\geqslant r$ in (3.1.20), we obtain
    $-sa_{2}+s\geqslant r+s\geqslant-1+a_{0}-t-a_{0}a_{1}+a_{0}a_{1}a_{2}\geqslant-1-a_{0}a_{1}+a_{0}a_{1}a_{2}.$ (3.1.21)
    Since $01-a_{2}<$ 0, then
    $s\leqslant\dfrac{-1-a_{0}a_{1}(1-a_{2})}{1-a_{2}}=\dfrac{1}{a_{2}-1}-a_{0}a_{1}\leqslant 1-a_{0}a_{1}.$
    Actually $s=1-a_{0}a_{1}.$ Indeed if $s=1-a_{0}a_{1}$ (which would force $a_{2}=2$) from (3.1.21) we can write
    $-s\geqslant r+s\geqslant-1+a_{0}-t+a_{0}a_{1}\geqslant-1+a_{0}a_{1}=-s.$
    In this inequality, we have that $t=a_{0}$. Thus, in the second inequality of (3.1.19) we obtain $s\leqslant-a_{0}a_{1}$, which contradicts the assumption $s=1-a_{0}a_{1}$. Therefore, we have $s=1-a_{0}a_{1}$ so we say
    $s\leqslant-a_{0}a_{1}.$ (3.1.22)
    Replacing (3.1.22) in (3.1.20), we obtain
    $r\geqslant-1+(a_{0}-t)+(-s-a_{0}a_{1})+a_{0}a_{1}a_{2}\geqslant-1+a_{0}a_{1}a_{2}.$ (3.1.23)
    On the other hand, from (3.1.13) and (3.1.22) we obtain
    $w_{4}+td\geqslant-sw_{1}\geqslant a_{0}a_{1}w_{1}.$
    Since $a_{1}w_{1}=d-w_{0}$ and $w_{4}=d-a_{0}w_{0}$ the previous inequality can be rewritten as
    $a_{0}(d-w_{0})\leqslant d-a_{0}w_{0}+td\Rightarrow a_{0}-1\leqslant t.$
    Also, from (3.1.19) we have $t\leqslant a_{0}$. Thus $t\in\{a_{0}-1,a_{0}\}$.
    Next, we will determine the solutions of the Diophantine equation (3.1.5) for each value of $t$.
    • i)
      If $t=a_{0}-1$, from (3.1.13) we obtain
      $w_{4}+(a_{0}-1)d+sd+rw_{2}\geqslant 0\Rightarrow w_{4}-d+a_{0}d+sd\geqslant-rw_{2}.$
      From (3.1.19) we have $r\leqslant-sa_{2}$. In addition, the weights verify $d=w_{4}+a_{0}w_{0}=w_{1}+a_{2}w_{2}$. Thus, the inequality above can be written
      $-a_{0}w_{0}+a_{0}d+sd\geqslant-rw_{2}\geqslant sa_{2}w_{2}=sd-sw_{1}.$
      As $a_{1}w_{1}=d-w_{0}$, then $-a_{0}a_{1}\leqslant s$. Also, by (3.1.22) we know that $s\leqslant-a_{0}a_{1}$. Hence $s=-a_{0}a_{1}$. On the other hand, replacing $t=a_{0}-1$ and $s=-a_{0}a_{1}$ in (3.1.20), we obtain $r\geqslant a_{0}a_{1}a_{2}$. From (3.1.19) we have
      $ra_{3}-w_{4}\leqslant t+s+r\Rightarrow r\leqslant\dfrac{w_{4}+t+s}{a_{3}-1}=\dfrac{w_{4}+(a_{0}-1)+(-a_{0}a_{1})}{a_{3}-1}.$
      As $w_{4}=1-a_{0}+a_{0}a_{1}-a_{0}a_{1}a_{2}+a_{0}a_{1}a_{2}a_{3}$, we obtain $r\leqslant a_{0}a_{1}a_{2}$. Thus, $r=a_{0}a_{1}a_{2}$. In this case, from (3.1.10) it follows that a solution for (3.1.5) is $(1,a_{1},0,0,0)$.
    • ii)
      If $t=a_{0}$. Following a similar process that in i) and using (3.1.19) and (3.1.22) we have
      $r\leqslant\dfrac{t+s+w_{4}}{a_{3}-1}=\dfrac{a_{0}+s+w_{4}}{a_{3}-1}\leqslant\dfrac{a_{0}-a_{0}a_{1}+w_{4}}{a_{3}-1}.$ (3.1.24)
      Moreover, since $w_{4}=1-a_{0}+a_{0}a_{1}-a_{0}a_{1}a_{2}+a_{0}a_{1}a_{2}a_{3}$, we obtain
      $r\leqslant\dfrac{a_{0}-a_{0}a_{1}+w_{4}}{a_{3}-1}\leqslant\dfrac{1-a_{0}a_{1}a_{2}+a_{0}a_{1}a_{2}a_{3}}{a_{3}-1}=\dfrac{1}{a_{3}-1}+a_{0}a_{1}a_{2}.$
      When $a_{3}=2$, it is possible that $r=1+a_{0}a_{1}a_{2}$. Let us see that this situation cannot happen. Indeed, if we assume that $r=1+a_{0}a_{1}a_{2}$, then $a_{3}=2$. It implies that $w_{4}=1-a_{0}+a_{0}a_{1}+a_{0}a_{1}a_{2}$. Replacing in (3.1.24), we obtain
      $r\leqslant 1+a_{0}a_{1}+s+a_{0}a_{1}a_{2}\leqslant 1+a_{0}a_{1}a_{2}=r.$
      Thus, we get $s=-a_{0}a_{1}$. Putting this in the third inequality of (3.1.19), we have $r\leqslant a_{0}a_{1}a_{2}$, which results in a contradiction. Therefore, we have $r\leqslant a_{0}a_{1}a_{2}$. Also, from (3.1.23), it verifies $r\geqslant-1+a_{0}a_{1}a_{2}$. Hence, we have $r\in\{-1+a_{0}a_{1}a_{2},a_{0}a_{1}a_{2}\}$. Next, we detail each case.
        If $r=-1+a_{0}a_{1}a_{2}$, we have in the last inequality of (3.1.19):
        $a_{0}+s-1+a_{0}a_{1}a_{2}=t+s+r\geqslant\dfrac{w_{3}-1}{a_{4}}=-1+a_{0}-a_{0}a_{1}+a_{0}a_{1}a_{2},$
        which implies that $s\geqslant-a_{0}a_{1}$. Using (3.1.22), we have $s=-a_{0}a_{1}$. Thus, in this case from (3.1.10) it follows that the solution of (3.1.5) is $(0,0,1,a_{3},0)$.
        If $r=a_{0}a_{1}a_{2}$, in a similar way as we have worked above, in the last inequality of (3.1.19) we have
        $a_{0}+s+a_{0}a_{1}a_{2}=t+s+r\geqslant\dfrac{w_{3}-1}{a_{4}}=-1+a_{0}-a_{0}a_{1}+a_{0}a_{1}a_{2},$
        which implies that $s\geqslant-1-a_{0}a_{1}$. From (3.1.22), $s\leqslant-a_{0}a_{1}$, we obtain $s=-1-a_{0}a_{1}$ or $s=-a_{0}a_{1}$. If $s=-1-a_{0}a_{1}$, the solution for (3.1.5) is $(0,1,a_{2},0,0)$. On the other hand, if $s=-a_{0}a_{1}$, from (3.1.10) it follows that $(0,0,0,1,a_{4})$ is the solution for the equation (3.1.5).
Therefore, the equation (3.1.5) has exactly five solutions:
$(a_{0},0,0,0,1),(1,a_{1},0,0,0),(0,1,a_{2},0,0),(0,0,1,a_{3},0)\mbox{ and }(0,0,0,1,a_{4}).$
Next, we compute the generators of $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$. For this, we need to find all monomials $z_{0}^{a_{0}}z_{1}^{a_{1}}z_{2}^{a_{2}}z_{3}^{a_{3}}z_{4}^{a_{4}}$ of degree $w_{i}$. It is equivalent to solve the Diophantine equation
$w_{0}x_{0}+w_{1}x_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=w_{i}$ (3.1.25)
with variables $x_{i}\in\mathbb{Z}_{0}^{+}$ and where at least one of them is nonzero.
Lemma 3.1.3.
The equation (3.1.25) has a unique solution for each $w_{i}$ and the set of generators of $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$ is given by $\{z_{i}\}$ for $i=0\ldots,4.$
Proof.
Since $f$ is a cycle polynomial, we can work without loss of generality with $w_{i}=w_{1}$. Thus, the equation (3.1.25) results
$w_{0}x_{0}+w_{1}x_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=w_{1}.$ (3.1.26)
From (3.1.2), we have $a_{2}w_{2}+w_{1}=d$. Then, if we add $a_{2}w_{2}$ to both sides of (3.1.26), we arrive to a new Diophantine equation
$w_{0}x_{0}+w_{1}x_{1}+w_{2}\tilde{x}_{2}+w_{3}x_{3}+w_{4}x_{4}=d,$ (3.1.27)
where the new variable $\tilde{x}_{2}=x_{2}+a_{2}\geqslant a_{2}$. Now, we remember that Equation (3.1.27) has five solutions, which were obtained in the lemma above:
$(a_{0},0,0,0,1),(1,a_{1},0,0,0),(0,1,a_{2},0,0),(0,0,1,a_{3},0)\mbox{ and }(0,0,0,1,a_{4}).$
Since $\tilde{x}_{2}\geqslant a_{2}$, we have a unique option that solve (3.1.27): $(0,1,a_{2},0,0)$. As a consequence, the equation (3.1.26) has a unique solution. Thus returning to Equation (3.1.26), after subtracting the vector $(0,0,a_{2},0,0)$ we obtain the solution $(0,1,0,0,0)$ that gives as generator the monomial $z_{1}.$ Similarly we obtain the other solutions which are exactly the solutions given by Equation (3.1.27). It follows that the generator for $H^{0}(\mathbb{P}(\mathbf{w}),w_{i})$ is given by the monomial $z_{i}$ for $i=0\ldots 4.$
Remark 3.1.1.
Let us choose the definition of the weighted projective space $\mathbb{P}(\mathbf{w})$ as a scheme $\operatorname{Proj}(S(\mathbf{w}))$, where
$S(\mathbf{w})=\bigoplus_{d}S^{d}(\mathbf{w})=\mathbb{C}\left[z_{0},z_{1},z_{2},z_{3},z_{4}\right].$
The ring of polynomials $\mathbb{C}\left[z_{0},z_{1},z_{2},z_{3},z_{4}\right]$ is graded with grading defined by the weights $\mathbf{w}=\left(w_{1},w_{1},w_{2},w_{3},w_{4}\right)$. From Lemma 3.1.3, the group $G(\mathbf{w})$ of automorphisms of the graded ring $S(\mathbf{w})$ can be defined on generators by
$\varphi_{\mathrm{w}}\left(\begin{array}[]{l}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\end{array}\right)=\left(\begin{array}[]{l}\alpha_{0}z_{0}\\ \alpha_{1}z_{1}\\ \alpha_{2}z_{2}\\ \alpha_{3}z_{3}\\ \alpha_{4}z_{4}\end{array}\right)$
where $\alpha_{i}\in\mathbb{C}^{*}.$ The group $\mathfrak{G}_{\mathbf{w}}$ of complex automorphisms of $\mathbb{P}(\mathbf{w})$ is the projectivization of $G(\mathbf{w})$ which in this case is given by $\mathfrak{G}_{\mathbf{w}}=\left(\mathbb{C}^{*}\right)^{4}.$ Actually since the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the following quasi-smooth monomials $z_{0}^{a_{0}}z_{4},z_{0}z_{1}^{a_{1}},z_{1}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}}.$ It follows (c.f. [6]) that the moduli of the orbifold $Z_{f},$ a $\mathbb{Q}$-homology projective 3-space, is given by
$Span\{z_{0}^{a_{0}}z_{4},z_{0}z_{1}^{a_{1}},z_{1}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}}\}/G(\mathbf{w})$
which determines a 0-dimensional quotient.
From the two previous lemmas, we obtain the following outcome in the context of Sasaki-Einstein structures for $\mathbb{Q}$-homology spheres.
Theorem 3.1.1.
Let $f$ be a cycle polynomial as in (3.1.1) of degree $d$ with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ such that $\gcd(d,w_{i})=1$ and index $I=|\textbf{w}|-d=1$. Then the complex dimension of the local moduli of the orbifold $Z_{f}=(f=0)/\mathbb{C}^{*}(\textbf{w})$, is equal to 00. Moreover, the generators of the spaces $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ and $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$ for all $i=0,\ldots,4$ are given in the previous lemmas of this subsection. Additionally, if $f$ belongs to one of the 236 $\mathbb{Q}$-homology spheres admitting Sasaki-Einstein metrics found in [5] and [13], then the dimension of the local moduli of Sasaki-Einstein metrics of $\mathbb{Q}$-homology 7-spheres at $L_{f}$ is zero dimensional. Thus $\mathbb{Q}$-homology 7-spheres given as links coming from polynomials $f$ as above do not admit inequivalent families of Sasaki-Einstein structures.
Proof.
The result follows from Lemmas 3.1.2, and 3.1.3. ∎
3.2. $\mathbb{Q}$-homology 7-spheres: Thom-Sebastiani sums of invertible polynomials
In this subsection we consider the following types of invertible polynomials:
Type I (Fermat-Cycle): $z_{0}^{a_{0}}+z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
Type II (Chain-Cycle): $z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
Type III (Cycle-Cycle): $z_{1}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}.$
From the Johnson and Kollar’s list [19] of well-formed anticanonically embedded Fano 3-folds of index 1, one can deduce that the weight vectors associated to these families of polynomials have the form:
$\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{3}v_{0},m_{3}v_{1},m_{2}v_{2},m_{2}v_{3},m_{2}v_{4}),$ (3.2.1)
where $\gcd(m_{2},m_{3})=1$ and, from well-formedness, $\gcd(v_{0},v_{1})=1$ and $\gcd(v_{i},v_{j})=1$ for $i\neq j$ with $i,j\in\{2,3,4\}$. Moreover, it is known that when a weight vector w as in (3.2.1) admits an invertible polynomial $f$ of degree $d=m_{2}m_{3}$ and index $I=|\textbf{w}|-d=1$, then the topology of its link $L_{f}$ correspond to a $\mathbb{Q}$-homology sphere, whose torsion can be calculated from the weights $w_{i}$’s.
Before we find the generator of the space of deformations of the orbifold $Z_{f}=\{f=0\}/\mathbb{C}^{*}(\textbf{w})$, we will prove some technical lemmas on the different types of polynomial described above.
Lemma 3.2.1.
Let $f$ be an invertible polynomial of type I, II or III with associated weight vector $\bf w$ described in (3.2.1) and with degree $d=m_{2}m_{3}$. Then we have
  • a)
    If $f$ is an invertible polynomial of type I, then $v_{0}=v_{1}=1$ in $\bf w$ and hence $w_{0}=w_{1}.$
  • b)
    If $f$ is an invertible polynomial of type II, then $v_{0}=1$ and if the pair $(\bf w,d)$ does not admit a polynomial of type I, then $v_{1}\neq 1$.
  • c)
    If $f$ is an invertible polynomial of type III, such that $(\bf w,d)$ does not admit a polynomial of type II, then $v_{0}\neq 1$ and $v_{1}\neq 1$.
Proof.
  • a)
    For a polynomial $f$ of type I:
    $f=z_{0}^{a_{0}}+z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
    with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ and degree $d=m_{2}m_{3}$, we have
    $\left\{\begin{array}[]{rcl}a_{0}w_{0}=d\Rightarrow a_{0}m_{3}v_{0}=m_{2}m_{3}\\ a_{1}w_{1}=d\Rightarrow a_{1}m_{3}v_{1}=m_{2}m_{3}.\end{array}\right.$
    The equalities above imply $v_{0}\mid m_{2}$ and $v_{1}\mid m_{2}$. Since $w_{i}=m_{2}v_{i}$ for $i=2,3,4$, we have $v_{0}\mid\gcd(w_{0},w_{1},w_{2},w_{3},w_{4})$. Finally, since w is well-formed, we get $v_{0}=1$. $v_{1}=1$ is obtained in a similar way.
  • b)
    For a polynomial $f$ of type II
    $f=z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
    with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ and degree $d=m_{2}m_{3}$, we have
    $a_{0}w_{0}=d\Rightarrow a_{0}m_{3}v_{0}=m_{2}m_{3}.$
    Thus $v_{0}\mid m_{2}$. As $w_{i}=m_{2}v_{i}$ for $i=2,3,4$, we also obtain $v_{0}\mid\gcd(w_{0},w_{2},w_{3},w_{4})$. Since w is well-formed, we conclude that $v_{0}=1$. Now, if we suppose that $v_{1}=1$, then we can choose the polynomial
    $\tilde{f}=z_{0}^{m_{2}}+z_{1}^{m_{2}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
    which is of type I for the weight vector $\bf w$, but this contradicts the hypothesis. Thus, we have $v_{1}\neq 1$.
  • c)
    We consider an invertible polynomial of type III:
    $f=z_{1}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}},$
    which has associated the weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ and degree $d=m_{2}m_{3}$. If we suppose that $v_{0}=1$, then $\bf w$ admits a polynomial $\tilde{f}$ of type II:
    $\tilde{f}=z_{0}^{m_{2}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$
    which is a contradiction. Thus, we have $v_{0}\neq 1$. A similar process as the one given above leads to $v_{1}\neq 1$.
In the next lemma, we show that the existence of certain type of polynomial for a given weight vector $\bf w$ implies the presence of a different type of invertible polynomial associated to the same weight vector $\bf w$.
Lemma 3.2.2.
Let $f$ be an invertible polynomial of type $I$, $II$ or $III$, where its associated weight vector $\bf w$ is described as in (3.2.1) and its degree satisfies $d=m_{2}m_{3}$. Then the following hold:
  • a)
    If $f$ is a polynomial of type II, then its associated weight vector $\bf w$ also admits a polynomial of type III.
  • b)
    If $f$ is a polynomial of type I, then its associated weight vector $\bf w$ also admits polynomials of type II and III.
Proof.
  • (a)
    Let $f$ be a polynomial of type II:
    $f=z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}.$
    By Lemma 3.2.1, we know $w_{0}=m_{3}$. As $d=m_{2}m_{3}$, we have $d-w_{1}=m_{3}(m_{2}-v_{1})>m_{3}$. This implies that $m_{3}\mid d-w_{1}$ and there exists an integer $\tilde{a}_{0}>1$ such that
    $\tilde{a}_{0}w_{0}=\tilde{a}_{0}m_{3}=d-w_{1}.$
    Thus, the weights vector $\bf w$ admits a polynomial $\tilde{f}$ of type III:
    $\tilde{f}=z_{1}z_{0}^{\tilde{a}_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}.$
  • (b)
    If $\bf w$ is the associated weight vector to the invertible polynomial $f$ of type I:
    $f=z_{0}^{a_{0}}+z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}},$
    then by Lemma 3.2.1, we have $w_{0}=w_{1}=m_{3}$. As $d=a_{1}m_{3}$ and $d>w_{0}+w_{1}=2m_{3}$, these imply that $a_{1}\geqslant 3$. Now, if we take $\tilde{a}_{1}=a_{1}-1\geqslant 2$, we can verify that $w_{0}+\tilde{a}_{1}w_{1}=d$. Therefore, $\bf w$ also admits a polynomial $\tilde{f}$ of type II:
    $\tilde{f}=z_{0}^{a_{0}}+z_{0}z_{1}^{\tilde{a}_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}.$
    Finally, from (a) of this lemma, we have that $\bf w$ also admits a polynomial of type III.
For any invertible polynomial $f$ of type I, II or III, whose associated weight vector $\bf w$ is defined as in (3.2.1) and with degree is $d=m_{2}m_{3}$, we have that $f$ contains no block of the form $z_{i}^{2}+z_{j}^{2}$ (recall $\gcd(m_{2},m_{3})=1$). This implies that $\dim\mathfrak{Aut}(Z_{f})=0$, see [7]. Thus, for the hypersurface $\{f=0\}\subset\mathbb{C}^{5}$, the complex dimension of the orbifold $Z_{f}\subset\mathbb{P}(\bf w)$ is given by the formula
$\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i})).$ (3.2.2)
We will begin computing the generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$. As before, we look for all different monomials $z_{0}^{y_{0}}z_{1}^{y_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $d$. Since $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ is the associated weight vector to $f$, equivalently we can solve the following Diophantine equation:
$w_{0}y_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d$ (3.2.3)
with variables $y_{i},x_{i}\in\mathbb{Z}^{+}_{0}$ and where at least one of them is nonzero. The following lemma allows us to split Equation (3.2.2) in two new Diophantine equations.
Lemma 3.2.3.
In the Diophantine equation (3.2.2), if some $y_{i}$ is a positive integer, then $x_{j}=0$, for all $j=2,3,4$. The converse is also true.
Proof.
On the contrary, let us assume that there exist some positive integers $y_{i}$ and $x_{j}$. In this case, we can suppose without loss of generality that $y_{0}>0$ and $x_{2}>0$. Since $\bf w$ is defined as in (3.2.1), we can write Equation (3.2.3) as:
$m_{3}v_{0}y_{0}+m_{3}v_{1}y_{1}+m_{2}v_{2}x_{2}+m_{2}v_{3}x_{3}+m_{2}v_{4}x_{4}=m_{2}m_{3}.$
Then, we have
$m_{2}(v_{2}x_{2}+v_{3}x_{3}+v_{4}x_{4})=m_{3}(m_{2}-v_{0}y_{0}-v_{1}y_{1}).$
Since $\gcd(m_{2},m_{3})=1$, then $m_{2}\mid(m_{2}-v_{0}y_{0}-v_{1}y_{1}).$ Moreover, since $m_{2}-v_{0}y_{0}-v_{1}y_{1}<m_{2}$, we obtain $m_{2}-v_{0}y_{0}-v_{1}y_{1}=0.$ Thus, we have
$m_{2}(v_{2}x_{2}+v_{3}x_{3}+v_{4}x_{4})=m_{3}(m_{2}-v_{0}y_{0}-v_{1}y_{1})=0.$
Finally, as $x_{2}>0$ and $x_{3},x_{4}\in\mathbb{Z}_{0}^{+}$, the equality above implies that $m_{2}=0$, which is not possible. For the converse, the process is similar to that in the previous case. ∎
From this lemma we conclude that the solutions of (3.2.3) can be obtained putting together the solutions of each one of the following Diophantine equations:
$w_{0}y_{0}+w_{1}y_{1}=d$ (3.2.4)
and
$w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d.$ (3.2.5)
Next, we solve Equations (3.2.4) and (3.2.5). The next lemma will allow us to find all the solutions of Equation (3.2.4).
Lemma 3.2.4.
Let $f$ be an invertible polynomial of type I, II or III, whose associated weight vector is $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfying (3.2.1) and with degree $d=m_{2}m_{3}$. Then we have:
  • a)
    If $f$ is of type I, then Equation (3.2.4) has $m_{2}+1$ solutions.
  • b)
    If $f$ is of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then Equation (3.2.4) has $\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}+1$ solutions.
  • c)
    If $f$ is of type III and its associated weight vector $\bf w$ does not admit polynomial of type II, then Equation (3.2.4) has $\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}+1$ solutions.
Moreover, all these quantities which represent the number of solutions of (3.2.4) are equivalent to
$\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor+1.$
Proof.
Indeed
  • a)
    Since $f$ is of type I, we have $v_{0}=v_{1}=1$. These imply that $w_{0}=w_{1}=m_{3}$. Replacing in Equation (3.2.4), we obtain
    $m_{3}y_{0}+m_{3}y_{1}=d=m_{2}m_{3}.$
    Simplifying, we obtain a new Diophantine equation $y_{0}+y_{1}=m_{2}$. Since all solutions are in $\mathbb{Z}_{0}^{+}$, these are given by the pairs
    $\{(0,m_{2}),(1,m_{2}-1),\dots(m_{2},0)\}.$
    Thus Equation (3.2.4) has $m_{2}+1$ solutions.
  • b)
    By Lemma 3.2.1, we have $v_{0}=1$ and $v_{1}\neq 1$. As a consequence, we obtain $w_{0}=m_{3}$. Moreover, as $d=m_{2}m_{3}$ and $w_{1}=m_{3}v_{1}$, the Equation (3.2.4) can be reduced to
    $y_{0}+v_{1}y_{1}=m_{2}.$ (3.2.6)
    Since $(m_{2},0)$ is a particular solution of (3.2.6), by Lemma 3.1.1 we have that all its solutions are given by the pairs
    $(y_{0},y_{1})=(m_{2}-kv_{1},k)\ \ \mbox{ where }k\in\mathbb{Z}.$
    Since $y_{0},y_{1}\in\mathbb{Z}_{0}^{+}$, we obtain
    $0\leqslant k\leqslant\dfrac{m_{2}}{v_{1}}.$ (3.2.7)
    Moreover, as $\bf w$ admits a polynomial of type II, we have $w_{0}+a_{1}w_{1}=d$. By above, it implies that $1+a_{1}v_{1}=m_{2}$. Thus, we conclude that $\dfrac{m_{2}-1}{v_{1}}$ is an integer. Considering this in (3.2.7), we write
    $0\leqslant k\leqslant\left(\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}\right)+\dfrac{1}{v_{1}}.$
    Moreover, as $v_{1}\neq 1$, we have the number of solutions of (3.2.4):
    $\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor+1=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}+1.$
  • c)
    Let $f$ be the polynomial of type III associated to $\bf w$:
    $f=z_{1}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}.$
    Replacing $w_{0}=m_{3}v_{0}$, $w_{1}=m_{3}v_{1}$ and $d=m_{2}m_{3}$ in (3.2.4) and then simplifying, we obtain the new equation
    $v_{0}y_{0}+v_{1}y_{1}=m_{2}$ (3.2.8)
    Since $\gcd(v_{0},v_{1})=1$ and the pair $(a_{0},1)$ is a solution of (3.2.8), we have that all solutions of the Diophantine equation (3.2.8) are given by the pairs
    $(y_{0},y_{1})=(a_{0}-kv_{1},1+kv_{0})$
    As $y_{0},y_{1}\in\mathbb{Z}_{0}^{+}$, then $k$ is restricted to
    $-\dfrac{1}{v_{0}}\leqslant k\leqslant\dfrac{a_{0}}{v_{1}}$ (3.2.9)
    Since $\bf w$ does not admit polynomials of type II, we have $v_{0}\neq 1$. Thus, we obtain $k\geqslant 0$. On the other hand, as the polynomial $f$ is associated to $\bf w$, we have $a_{0}w_{0}+w_{1}=d$. This implies that $a_{0}v_{0}+v_{1}=m_{2}$. Then
    $\dfrac{a_{0}}{v_{1}}=\dfrac{m_{2}-v_{1}}{v_{0}v_{1}}=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}=\left(\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}\right)+\dfrac{1}{v_{1}}.$
    Using again that $\bf w$ does not admit polynomial of type II, we have $v_{1}\neq 1$, which implies that $\dfrac{1}{v_{1}}<1$. In addition, since the link $L_{f}$ is a $\mathbb{Q}$-homology sphere, we have that $\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}$ is a integer. Therefore, in (3.2.9) we have:
    $0\leqslant k\leqslant\left\lfloor\dfrac{a_{0}}{v_{1}}\right\rfloor=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}$
    Finally, as $\gcd(v_{0},v_{1})=1$, then $0<\dfrac{1}{v_{0}}+\dfrac{1}{v_{1}}$. From $\gcd(v_{0},v_{1})=1,$ it follows that the number of solutions of (3.2.4) is
    $\left\lfloor\dfrac{a_{0}}{v_{1}}\right\rfloor+1=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}+1=\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor+1.$
Remark 3.2.1.
We have the following remarks.
  • a)
    Notice that if $f$ is a polynomial of type II such that its associated weight vector $\bf w$ admits a polynomial of type I, then we can work as in a) of lemma above. Therefore, the equation (3.2.4) has $m_{2}+1$ solutions.
  • b)
    On the other hand, if $f$ is a polynomial of type III whose associated weight vector $\bf w$ admits a polynomial of type II but not type I, then we can use the case b) of the previous lemma.
  • c)
    Finally, if $f$ is a polynomial of type III whose associated weight vector admits polynomial of type I and II, then the number of solutions of (3.2.4) is obtained as in a) of the previous lemma.
It remains to compute the number of solutions of Equation (3.2.5). Due to the fact that the polynomials of type I, II or III have the same cycle block:
$z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}},$ (3.2.10)
the number of solutions of the Diophantine Equation (3.2.5) is independent of any type. On the other hand, as the polynomial $f$ has degree $d=m_{2}m_{3}$ and $\bf w$ is written as in (3.2.1), then we obtain from (3.2.10) the following equations for the weights $w_{i}$’s:
$a_{2}w_{2}+w_{4}=d$ $\Rightarrow v_{4}+a_{2}v_{2}=m_{3}$ (3.2.11)
$a_{3}w_{3}+w_{2}=d$ $\Rightarrow v_{2}+a_{3}v_{3}=m_{3}$ (3.2.12)
$a_{4}w_{4}+w_{3}=d$ $\Rightarrow v_{3}+a_{4}v_{4}=m_{3}.$ (3.2.13)
In addition, as the link $L_{f}$ is a $\mathbb{Q}$-homology sphere, it verifies Equality (2.1.4): $m_{3}=a_{2}a_{3}a_{4}+1$. Thus, we can write $v_{2}$, $v_{3}$ and $v_{4}$ as
$v_{2}=a_{4}a_{3}-a_{3}+1,\ \ v_{3}=a_{2}a_{4}-a_{4}+1,\ \ \mbox{ and }\ v_{4}=a_{3}a_{2}-a_{2}+1.$ (3.2.14)
Next, we will solve Equation (3.2.5):
Lemma 3.2.5.
The Diophantine equation
$w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d,$
where $w_{i}$’s are defined as in (3.2.1), has only three solutions.
Proof.
Since $w_{j}=m_{2}v_{j}$, for $j=2,3,4$ and $d=m_{2}m_{3}$, we have an equivalent equation to (3.2.5):
$v_{2}x_{2}+v_{3}x_{3}+v_{4}x_{4}=m_{3}.$ (3.2.15)
As $\gcd(v_{2},v_{3})=1$, we can define a new variable $\tilde{x}=v_{2}x_{2}+v_{3}x_{3}.$ Thus, we can write Equation (3.2.15) as
$\tilde{x}+v_{4}x_{4}=m_{3}.$ (3.2.16)
By (3.2.13), a particular solution of the equation (3.2.16) is given by the pair $(v_{3},a_{4}).$ Then the general solution of (3.2.16) is
$(\tilde{x},x_{4})=(v_{3}+tv_{4},a_{4}-t),\ \ \mbox{ where }t\in\mathbb{Z}.$ (3.2.17)
Now, we consider the Diophantine equation
$v_{2}x_{2}+v_{3}x_{3}=\tilde{x}=v_{3}+tv_{4}$ (3.2.18)
From (3.2.11) and (3.2.12), we have
$v_{4}=m_{3}-a_{2}v_{2}=(v_{2}+a_{3}v_{3})-a_{2}v_{2}=a_{3}v_{3}+(1-a_{2})v_{2}.$
Therefore, a particular solution of (3.2.18) is the pair $(t(1-a_{2}),1+ta_{3})$. Then the general solution of (3.2.18) is given by
$(x_{2},x_{3})=(t(1-a_{2})+sv_{3},1+ta_{3}-sv_{2}),\ \ \mbox{ where }t,s\in\mathbb{Z}.$ (3.2.19)
From (3.2.17) and (3.2.19), we obtain the general solution of the Diophantine equation (3.2.15):
$(x_{2},x_{3},x_{4})=(t(1-a_{2})+sv_{3},1+ta_{3}-sv_{2},a_{4}-t),\ \ \mbox{ where }t,s\in\mathbb{Z}.$ (3.2.20)
Claim: $\bf s\in\{0,1\}$
Since we require $x_{2},x_{3}\in\mathbb{Z}_{0}^{+}$, we have that $x_{2}=t(1-a_{2})+sv_{3}\geqslant 0$ and $x_{3}=1+ta_{3}-sv_{2}\geqslant 0$. This inequalities imply that
$\dfrac{sv_{2}-1}{a_{3}}\leqslant t\leqslant\dfrac{sv_{3}}{a_{2}-1}.$
From this, we obtain
$\dfrac{sv_{2}-1}{a_{3}}\leqslant\dfrac{sv_{3}}{a_{2}-1}$ $\Rightarrow(a_{2}-1)(sv_{2}-1)\leqslant sv_{3}a_{3}$
$\Rightarrow-(a_{2}-1)\leqslant s(v_{3}a_{3}-v_{2}(a_{2}-1)).$
Using (3.2.11) and (3.2.12) we have $v_{4}=v_{3}a_{3}-v_{2}(a_{2}-1).$ Replacing above, we can write
$-(a_{2}-1)\leqslant sv_{4}\Rightarrow-\dfrac{a_{2}-1}{v_{4}}\leqslant s.$
Now, using the expression for $v_{4}$ given in (3.2.14) and the fact of that $a_{2}a_{3}\geqslant 2a_{2}>2a_{2}-1,$ we have
$v_{4}=a_{2}a_{3}-a_{2}+1>a_{2}\Rightarrow\dfrac{a_{2}-1}{v_{4}}<1.$
Hence
$-1<-\dfrac{a_{2}-1}{v_{4}}\leqslant s.$ (3.2.21)
On the other hand, as $x_{3}\geqslant 0$ and $x_{4}\geqslant 0$, then $1+ta_{3}\geqslant sv_{2}$ and $a_{4}\geqslant t$, respectively. Putting together these two inequalities, we obtain
$1+a_{4}a_{3}\geqslant 1+ta_{3}\geqslant sv_{2}.$
Also, from the expression given for $v_{2}$ given in (3.2.14) and the inequality $a_{4}a_{3}\geqslant 2a_{3}>2a_{3}-1$, we have
$2v_{2}=2a_{4}a_{3}-2a_{3}+2>1+a_{4}a_{3}.$
Since $sv_{2}\leqslant 1+a_{4}a_{3}$, we conclude that
$s\leqslant\dfrac{1+a_{4}a_{3}}{v_{2}}<2.$ (3.2.22)
From (3.2.21) and (3.2.22), we conclude that $s\in\{0,1\}.$
Next, we will exhibit the solutions that are obtained for each $s$.
For $\bf s=0$: In this case, the general solution of (3.2.15) is given by
$(x_{2},x_{3},x_{4})=(t(1-a_{2}),1+ta_{3},a_{4}-t).$
Since $x_{i}\geqslant 0$, we have $t(1-a_{2})\geqslant 0$, $1+ta_{3}\geqslant 0$ and $a_{4}-t\geqslant 0$. As $a_{2}-1>0$, then $t\leqslant 0$. On the other hand, the inequality $1+ta_{3}\geqslant 0$ implies that $t\geqslant-\frac{1}{a_{3}}>-1$ so $t=0$. Therefore, $(0,1,a_{4})$ is the unique solution for Equation (3.2.15).
For $\bf s=1$: The general solution of (3.2.15) is given by
$(x_{2},x_{3},x_{4})=(t(1-a_{2})+v_{3},1+ta_{3}-v_{2},a_{4}-t).$
Since $x_{4}=a_{4}-t\geqslant 0$, we have $a_{4}\geqslant t$. On the other hand, as $x_{3}=1+ta_{3}-v_{2}\geqslant 0$ and $v_{2}=a_{4}a_{3}-a_{3}+1$ in (3.2.14), we obtain
$t\geqslant\dfrac{v_{2}-1}{a_{3}}=a_{4}-1$
Thus, $t\in\{a_{4}-1,a_{4}\}$. So the two solutions for $s=1$ are
$\left\{\begin{array}[]{cl}(1,a_{3},0),&\hbox{ if }t=a_{4},\\ (a_{2},0,1),&\hbox{ if }t=a_{4}-1.\end{array}\right.$
Remark 3.2.2.
From the previous two lemmas we have that for $f$ be an invertible polynomial with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfying (3.2.1) and with degree $d=m_{2}m_{3}$.
  • If $f$ is a polynomial of type I, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w},\mathcal{O}(d))$ is given by the following monomials of degree $d$:
    $\{z_{1}^{m_{2}},z_{0}z_{1}^{m_{2}-1},\dots,z_{0}^{m_{2}-1}z_{1},z_{0}^{m_{2}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}}\}.$
  • If $f$ is a polynomial of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w},\mathcal{O}(d))$ is given by the following monomials of degree $d:$
    $\left\{z_{0}^{m_{2}-kv_{1}}z_{1}^{k},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor+1\right\}.$
  • If $f$ is a polynomial of type III and its associated weight vector $\bf w$ does not admit polynomial of type II,, then the set of generators of $H^{0}(\mathbb{P}(\textbf{w},\mathcal{O}(d))$ is given by the following monomials of degree $d:$
    $\left\{z_{0}^{a_{0}-kv_{1}}z_{1}^{1+kv_{0}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor+1\right\}.$
Now we will compute $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))$ for each $w_{i}$, that is, we will find all the solutions of the following Diophantine equation
$w_{0}y_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=w_{i},$ (3.2.23)
with $x_{j},y_{k}\in\mathbb{Z}_{0}^{+}$, where at least one of them is nonzero. In the next lemma, we will do this for either $w_{i}=w_{0}$ or $w_{i}=w_{1}.$
Lemma 3.2.6.
Let $f$ be an invertible polynomial of type I, II or III, with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfying (3.2.1) and with degree $d=m_{2}m_{3}$. We consider the following Diophantine equations
$w_{0}y_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}$ $=w_{0}$ (3.2.24)
$w_{0}y_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}$ $=w_{1}.$ (3.2.25)
Then, we have:
  • a)
    If $f$ is of type I, then equations (3.2.24) and (3.2.25) both have two solutions. Moreover, the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))$ is given by $\{z_{0},z_{1}\}$ for $i=0,1.$
  • b)
    If $f$ is of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then equations (3.2.24) and (3.2.25) have one and two solutions, respectively. Moreover, the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{0}))$ is given by $\{z_{0}\}$ and the set generators for $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{1}))$ is $\{z_{0}^{v_{1}},z_{1}\}.$
  • c)
    If $f$ is of type III and its associated weight vector $\bf w$ does not admit polynomial of type II, then the equations (3.2.24) and (3.2.25) both have one solution. Moreover, the set of generators for $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{0}))$ is $\{z_{0}\}$ and the set of generators for $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{1}))$ is $\{z_{1}\}$
Proof.
Let us study each case:
  • (a)
    If $f$ is of type I, we can write
    $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{3},m_{3},m_{2}v_{2},m_{2}v_{3},m_{2}v_{4}).$
    Since $w_{0}=m_{3}$ and $d=m_{2}m_{3}$, adding $(m_{2}-1)w_{0}$ on both sides of (3.2.24), we obtain
    $w_{0}\hat{y}_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d,$ (3.2.26)
    where the new variable $\hat{y}_{0}=y_{0}+(m_{2}-1)\geqslant m_{2}-1$. From Lemmas 3.2.4 and 3.2.5, it follows that the solutions of Equation (3.2.26) belong to the set
    $\,\qquad\{(t,m_{2}-t,0,0,0),(0,0,a_{2},0,1),(0,0,1,a_{3},0),(0,0,0,1,a_{4}),\mbox{ where }t\in\mathbb{Z}_{0}^{+},t\leqslant m_{2}\}.$
    Since $\hat{y}_{0}\geqslant m_{2}-1$, we will find only two possible solutions for (3.2.26): if $\hat{y}_{0}=m_{2}$, we have $y_{0}=1$ and $y_{1}=0$. On the other hand, if $\hat{y}_{0}=m_{2}-1$, we obtain $y_{0}=0$ and $y_{1}=1$. Thus, the solutions obtained are
[…]
    Since $w_{1}=w_{0}$, we also obtain two solutions for Equation (3.2.25).
  • (b)
    First, we will solve Equation (3.2.24). Since $f=z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}}$ is a polynomial of type II and its associated weight vector $\bf w$ does not admit polynomials of type I, we can express $\bf w$ as
    $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{3},m_{3}v_{1},m_{2}v_{2},m_{2}v_{3},m_{2}v_{4}),\ \ \mbox{ where }v_{1}\neq 1.$
    As $w_{0}=m_{3}$ and $d=m_{2}m_{3}$, then adding $(m_{2}-1)w_{0}$ on both sides of the Equation (3.2.24), we obtain
    $w_{0}\hat{y}_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d$ (3.2.27)
    with new variable $\hat{y}_{0}=y_{0}+(m_{2}-1)\geqslant m_{2}-1$. By Lemmas 3.2.4 and 3.2.5, the solutions of (3.2.27) belong to the set
    $\qquad\left\{(m_{2}-tv_{1},t,0,0,0),(0,0,a_{2},0,1),(0,0,1,a_{3},0),(0,0,0,1,a_{4}),\mbox{ where }t\in\mathbb{Z}_{0}^{+},t\leqslant\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor\right\}$
    Notice that this is possible only when $\hat{y}_{0}=m_{2}-tv_{1}\geqslant m_{2}-1$. Since $v_{1}\neq 1$ then $t\leqslant\dfrac{1}{v_{1}}<1$. Thus, we conclude that $t=0$. In this case, we obtain $\hat{y}_{0}=m_{2}$, which implies $y_{0}=1$ and $y_{1}=0$. Hence, the unique solution for (3.2.24) is $(1,0,0,0,0).$
    On the other hand, to solve Equation (3.2.25), consider the term $z_{0}z_{1}^{a_{1}}$ of $f$, that leads to $w_{0}+a_{1}w_{1}=d.$ If we add $w_{0}+(a_{1}-1)w_{1}$ to both sides of Equation (3.2.25), we obtain the new Diophantine equation
    $w_{0}\hat{y}_{0}+w_{1}\hat{y}_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d,$ (3.2.28)
    where $\hat{y}_{0}=y_{0}+1\geqslant 1$ and $\hat{y}_{1}=y_{1}+a_{1}-1\geqslant a_{1}-1$. By the of solutions given above, we can write $\hat{y}_{0}=m_{2}-tv_{1}$ and $\hat{y}_{1}=t$. Thus, we obtain
    $\dfrac{m_{2}-1}{v_{1}}-1=a_{1}-1\leqslant t\leqslant\dfrac{m_{2}-1}{v_{1}}=\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor.$
    We have
    • If $t=\dfrac{m_{2}-1}{v_{1}}-1$, then $\hat{y}_{0}=1+v_{1}$ and $\hat{y}_{1}=\dfrac{m_{2}-1}{v_{1}}-1$. Thus, $y_{0}=v_{1}$ and $y_{1}=0$.
    • If $t=\dfrac{m_{2}-1}{v_{1}}$, then $\hat{y}_{0}=1$ and $\hat{y}_{1}=\dfrac{m_{2}-1}{v_{1}}$. Thus, $y_{0}=0$ and $y_{1}=1$.
    Thus, Equation (3.2.25) has two solutions.
  • (c)
    We consider $f$ be an invertible polynomial of type III:
    $f=z_{1}z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}},$
    whose associated weight vector $\bf w$ does not admit polynomial of type II. To solve (3.2.24), we add $a_{1}w_{1}$ to both sides of this equation. Since $w_{0}+a_{1}w_{1}=d$, we obtain
    $w_{0}y_{0}+w_{1}\hat{y}_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d$ (3.2.29)
    where $\hat{y}_{1}=y_{1}+a_{1}\geqslant a_{1}$. By Lemmas 3.2.4 c) and 3.2.5, the solutions of (3.2.29) are in the set
    $\qquad\left\{(a_{0}-tv_{1},1+tv_{0},0,0,0),(0,0,a_{2},0,1),(0,0,1,a_{3},0),(0,0,0,1,a_{4}),\mbox{ where }t\in\mathbb{Z}_{0}^{+},t\leqslant\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor\right\}$
    Then $\hat{y}_{1}=1+tv_{0}\geqslant a_{1}$. Thus $t\geqslant\dfrac{a_{1}-1}{v_{0}}$. As $w_{0}+a_{1}w_{1}=d$, we have that $v_{0}+a_{1}v_{1}=m_{2}$. We obtain
    $t\geqslant\dfrac{a_{1}-1}{v_{0}}=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}=\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor.$
    So $t=\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor$. Thus, Equation (3.2.29) has only one solution given by $(a_{0}-tv_{1},1+tv_{0},0,0,0).$ Returning to Equation (3.2.24) (after subtracting the vector $(0,a_{1},0,0,0)$) we obtain the vector $(a_{0}-tv_{1},1+tv_{0}-a_{1},0,0,0)$ as solution. In this case, since $t=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}},$ it follows that the generator of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{0}))$ is given by $\{z_{0}\}.$ Solutions to Equation (3.2.25) can be found using a similar argument as the one given above. It follows that the generator of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{1}))$ is given by $\{z_{1}\}.$
Remark 3.2.3.
We have the following comments.
  • a)
    Notice that if $f$ is a polynomial of type II such that its associated weight vector $\bf w$ admits a polynomial of type I, then part a) of the previous lemma above holds. Thus, Equations (3.2.24) and (3.2.25) have $2$ solutions each one.
  • b)
    If $f$ is a polynomial of type III whose associated weight vector $\bf w$ admits a polynomial of type II but not type I, then we can use part b) of the previous lemma above. Thus, Equations (3.2.24) and (3.2.25) have $1$ and $2$ solutions, respectively.
  • c)
    Finally, if $f$ is a polynomial of type III whose associated weight vector admits polynomial of type I and II, then the number of solutions of the equations can be obtained as in a) of the previous lemma. That is, Equations (3.2.24) and (3.2.25) both have $2$ solutions.
Next, we will solve Equation (3.2.23) for the remaining cases: $i=2,3$ or $4$.
Lemma 3.2.7.
Let $f$ be a polynomial of type I, II or III and $\bf w$ the associated weight vector defined as in (3.2.1). Then the equation (3.2.23):
$w_{0}y_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=w_{i}$
has a unique solution for each $i=2,3,4$. It follows that the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))$ is given by $\{z_{i}\}$ for $i=2,3,4.$
Proof.
Since any polynomial $f$ of type I, II or III has the same block of cycle type
$z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}},$
we can work with any of these types. On the other hand, by the cyclic form of this block, it is enough to assume $w_{i}=w_{2}$. Thus, Equation (3.2.23) can be written as
$w_{0}y_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=w_{2}$ (3.2.30)
Adding $a_{3}w_{3}$ to both sides of (3.2.30), we obtain
$w_{0}y_{0}+w_{1}y_{1}+w_{2}x_{2}+w_{3}\hat{x}_{3}+w_{4}x_{4}=d$
where $\hat{x}_{3}=x_{3}+a_{3}\geqslant a_{3}$. Since $\hat{x}_{3}>0$, the solutions of the above equation are in the set
$\left\{(0,0,a_{2},0,1),(0,0,1,a_{3},0),(0,0,0,1,a_{4})\right\}.$
Since $a_{3}>1$, we have only one solution $(0,0,1,a_{3},0)$. From $\hat{x}_{3}=a_{3}$, we obtain $x_{3}=0$. Then the solution of Equation (3.2.30) is $(0,0,1,0,0)$. Similarly we obtain the solutions $(0,0,0,1,0)$ and (0,0,,0,1)(0,0,,0,1)( 0 , 0 , , 0 , 1 ) for $w_{3}$ and $w_{4}$ respectively. The last statement of the lemma follows for similar arguments as the ones given in the previous lemmas. ∎
In the same vein as Remark 3.1.1 we give the next remark on the moduli of the orbifold associated to the polynomials considered in this subsection.
Remark 3.2.4.
The group $G(\mathbf{w})$ of complex automorphisms of the graded ring $S(\mathbf{w})$ with $\operatorname{Proj}(S(\mathbf{w}))$ can be defined on generators for polynomials of type I, II and III thanks to Lemmas 3.2.6 and 3.2.7. Notice that the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ in all these cases is given by quasi-smooth monomials, so we can describe the local moduli of the corresponding orbifold, a $\mathbb{Q}$-homology projective 3-space, as before. Indeed, let $\alpha_{i},\beta_{1}\in\mathbb{C}^{*}$ and $\mathbb{A}\in GL(2,\mathbb{C}),$ then we have
  • If $f$ is a polynomial of type I, then $G(\mathbf{w})$ is given on generators by
    $\varphi_{\mathrm{w}}\left(\begin{array}[]{l}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\end{array}\right)=\left(\begin{array}[]{c}\mathbb{A}\left(\begin{array}[]{l}z_{0}\\ z_{1}\end{array}\right)\\ \alpha_{2}z_{2}\\ \alpha_{3}z_{3}\\ \alpha_{4}z_{4}\end{array}\right)$
    and the moduli of the orbifold $Z_{f}$ can be expressed by the quotient
    $Span\left\{z_{1}^{m_{2}},z_{0}z_{1}^{m_{2}-1},\ldots,z_{0}^{m_{2}-1}z_{1},z_{0}^{m_{2}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}}\right\}/G(\mathbf{w}).$
  • If $f$ is a polynomial of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then $G(\mathbf{w})$ is given on generators by
    $\varphi_{\mathrm{w}}\left(\begin{array}[]{l}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\end{array}\right)=\left(\begin{array}[]{c}\alpha_{0}z_{0}\\ \alpha_{1}z_{1}+\beta_{1}z_{0}^{v_{1}}\\ \alpha_{2}z_{2}\\ \alpha_{3}z_{3}\\ \alpha_{4}z_{4}\end{array}\right)$
    and the moduli of the orbifold $Z_{f}$ can be expressed by the quotient
    $Span\left\{z_{0}^{m_{2}-kv_{1}}z_{1}^{k},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{1}}\right\rfloor+1\right\}/G(\mathbf{w}).$
  • Finally, If $f$ is a polynomial of type III and its associated weight vector $\bf w$ does not admit polynomial of type II, then $G(\mathbf{w})$ is given on generators by
    $\varphi_{\mathrm{w}}\left(\begin{array}[]{l}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\end{array}\right)=\left(\begin{array}[]{c}\alpha_{0}z_{0}\\ \alpha_{1}z_{1}\\ \alpha_{2}z_{2}\\ \alpha_{3}z_{3}\\ \alpha_{4}z_{4}\end{array}\right)$
    and the moduli of the orbifold $Z_{f}$ can be expressed by the quotient
    $Span\left\{z_{0}^{a_{0}-kv_{1}}z_{1}^{1+kv_{0}},z_{4}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}},\mbox{ where }0\leqslant k\leqslant\left\lfloor\dfrac{m_{2}}{v_{0}v_{1}}\right\rfloor+1\right\}/G(\mathbf{w}).$
For the local moduli of Sasaki-Einstein structures for $\mathbb{Q}$-homology spheres we can say that this is non-trivial, more precisely we have:
Theorem 3.2.1.
Let $f$ be an invertible polynomial of type I, II or III, whose associated weight vector is $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ which is defined as in (3.2.1) with degree $d=m_{2}m_{3}$. Then the complex dimension $\mu$ of the local moduli of the orbifold $Z_{f}=\{f=0\}/\mathbb{C}^{*}(\textbf{w})$ is given as follows
  • a)
    If $f$ is of type I, then $\mu=m_{2}-3$.
  • b)
    If $f$ is of type II and its associated weight vector $\bf w$ does not admit polynomial of type I, then $\mu=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2$
  • c)
    If $f$ is of type III and its associated weight vector $\bf w$ does not admit polynomial of type II, then $\mu=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1$.
All these quantities are equivalent to
$\mu=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$
Moreover, the generators of the spaces $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ and $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$ for all $i=0,\ldots,4$ are given in the previous lemmas of this subsection. Additionally, if $f$ belongs to one of the 236 $\mathbb{Q}$-homology spheres admitting Sasaki-Einstein metrics found in [5] and [13], then the real dimension of the local moduli of Sasaki-Einstein metrics of $\mathbb{Q}$-homology 7-spheres at $L_{f}$ equals $2\left[\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1\right]$.
Proof.
Let us prove these statements case by case:
  • a)
    If $f$ is of type I, from Lemmas 3.2.4 and 3.2.5 we know $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))=m_{2}+4$. Also, from Lemmas 3.2.6 and 3.2.7, we have
    $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=7.$
    Thus
    $\mu=\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=m_{2}-3.$
  • b)
    If $f$ is of type II with associated weight vector $\bf w$ not admitting a polynomial of type I, from Lemmas 3.2.4 and 3.2.5 we know $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}+4.$ On the other hand, Lemmas 3.2.6 and 3.2.7, establish the equality
    $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=6.$
    Thus
    $\mu=\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2.$
  • c)
    If $f$ is of type III and its associated weight vector $\bf w$ does not admit a polynomial of type II, then we obtain from Lemmas 3.2.4 and 3.2.5 that
    $\dim H^{0}(\mathbb{P}(\textbf{w}),d)=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}+4.$
    Moreover, from Lemmas 3.2.6 and 3.2.7, we have
    $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(w_{i}))=5.$
    Thus,
    $\mu=\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$
Aso we notice that for any of the cases mentioned above, $\mu$ is equivalent to
$\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$
The last statement then follows from Theorem 2.2.2
Remark 3.2.5.
We have some comments.
  • a)
    When $f$ is a polynomial of type II such that its associated weight vector $\bf w$ admits a polynomial of type I, then part a) of the previous theorem holds. Thus, $\mu=m_{2}-3$.
  • b)
    If $f$ is a polynomial of type III whose associated weight vector $\bf w$ admits a polynomial of type II but not type I, then we can use part b) of the previous theorem. Thus, $\mu=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2.$
  • c)
    Finally, if $f$ is a polynomial of type III whose associated weight vector admits polynomial of type I and II, then $\mu=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$
3.3. $\mathbb{Q}$-homology sphere: the Berglund-Hübsch dual of chain-cycle polynomials
As mentioned before, in [14] it was proven that the Berglund-Hübsch transpose rule only produces twins for singularities of cycle type or of type I and type III. Actually the types are preserved under the Berglund-Hübsch transpose:
  • Type I polynomials are sent to type I polynomials.
  • Type III polynomials are sent to type III polynomials, and moreover
  • $m_{2}$, $v_{0}$ and $v_{1}$ are invariant under the Berglund-Hübsch rule.
From Theorem 3.2.1 the real dimension $\mu_{\mathbb{R}}$ of the local moduli for links arising from polynomials of type I and type III are given by $\mu_{\mathbb{R}}=2[\frac{m_{2}}{v_{0}v_{1}}-\frac{1}{v_{0}}-\frac{1}{v_{1}}-1]$, and from Theorem 3.1.1 $\mu_{\mathbb{R}}=0$ in case the singularity is given by a cycle type polynomial given in Subsection 3.1. Thus we only need to study polynomials of type II, that is polynomials of chain-cycle type, where the Berglund-Hübsch transpose rule does not preserve neither torsion nor Milnor number, and hence $m_{2}$, $v_{0}$ and $v_{1}$ vary.
Let
$z_{0}^{a_{0}}+z_{0}z_{1}^{a_{1}}+z_{4}z_{2}^{a_{2}}+z_{2}z_{3}^{a_{3}}+z_{3}z_{4}^{a_{4}},$ (3.3.1)
where its associated weight vector $\bf w$ satisfies condition (3.2.1). The exponential matrix of $f$ is given by
$A_{f}=\begin{bmatrix}a_{0}&0&0&0&0\\ 1&a_{1}&0&0&0\\ 0&0&a_{2}&0&1\\ 0&0&1&a_{3}&0\\ 0&0&0&1&a_{4}\end{bmatrix}$ (3.3.2)
Applying the Berglund-Hübsch transpose rule, one obtains the matrix
$A^{T}_{f}=\begin{bmatrix}a_{0}&1&0&0&0\\ 0&a_{1}&0&0&0\\ 0&0&a_{2}&1&0\\ 0&0&0&a_{3}&1\\ 0&0&1&0&a_{4}\end{bmatrix}$
which has associated invertible polynomial
$f^{T}=z_{0}^{a_{0}}z_{1}+z_{1}^{a_{1}}+z_{3}z_{2}^{a_{2}}+z_{4}z_{3}^{a_{3}}+z_{2}z_{4}^{a_{4}},$ (3.3.3)
In [14], the associated weight vector $\tilde{\bf w}$ to $f^{T}$ was obtained:
$\tilde{\bf w}=(m_{3}v_{1}(a_{1}-1),m_{3}m_{2}v_{1},m_{2}(m_{2}-1)\tilde{v}_{2},m_{2}(m_{2}-1)\tilde{v}_{3},m_{2}(m_{2}-1)\tilde{v}_{4}),$
where the degree is $\tilde{d}=m_{3}m_{2}(m_{2}-1)$ and
$\tilde{v}_{2}=a_{3}a_{4}-a_{4}+1,\ \ \ \tilde{v}_{3}=a_{2}a_{4}-a_{2}+1\ \ \mbox{ and }\ \ \tilde{v}_{4}=a_{2}a_{3}-a_{3}+1.$ (3.3.4)
Also, since $m_{2}-1=a_{1}v_{1}$, we can simplify $\tilde{\bf w}$ and obtain
$\tilde{\bf w}=(m_{3}(a_{1}-1),m_{3}m_{2},m_{2}a_{1}\tilde{v}_{2},m_{2}a_{1}\tilde{v}_{3},m_{2}a_{1}\tilde{v}_{4}),$ (3.3.5)
where the degree is $\tilde{d}=m_{3}m_{2}a_{1}$.
The complex dimension of the local moduli for $Z_{f^{T}}$ is given by
$\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))-\sum_{i}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ (3.3.6)
We begin computing $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))$. Since this refers to the number of all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $\tilde{d}$, we calculate this number counting the solutions of the Diophantine equation
$\tilde{w}_{0}x_{0}+\tilde{w}_{1}x_{1}+\tilde{w}_{2}x_{2}+\tilde{w}_{3}x_{3}+\tilde{w}_{4}x_{4}=\tilde{d}$ (3.3.7)
where the unknowns $x_{i}$ are non-negative integers such that at least one of them is non-zero. We will solve this equation adding certain mild constraints on the exponent $a_{1}$ and $m_{3}$. As we will see the remaining cases of interest, those whose corresponding links admit Sasaki-Einstein metrics, can be computed case by case.
Lemma 3.3.1.
Let $f^{T}$ be an invertible polynomial described as in (3.3.3) with an associated weight vector $\tilde{\bf w}$ given as in (3.3.5). If we add the additional conditions $\gcd(a_{1},m_{3})=1$ and $a_{1}>2$, then the equation (3.3.7) has exactly five solutions. Moreover, the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(\tilde{d}))$ is given by the set
$\{z_{0}^{m_{2}}z_{1},z_{1}^{a_{1}},z_{3}z_{2}^{a_{2}},{z_{4}}z_{3}^{a_{3}},z_{2}z_{4}^{a_{4}}\}.$
Proof.
Using the expression given in (3.3.5), we can write the Diophantine equation (3.3.7) as
$m_{3}(a_{1}-1)x_{0}+m_{3}m_{2}x_{1}+m_{2}a_{1}\tilde{v}_{2}x_{2}+m_{2}a_{1}\tilde{v}_{3}x_{3}+m_{2}a_{1}\tilde{v}_{4}x_{4}=m_{3}m_{2}a_{1},$ (3.3.8)
which we can rewrite as
$m_{3}((a_{1}-1)x_{0}+m_{2}x_{1})=a_{1}m_{2}(m_{3}-\tilde{v}_{2}x_{2}-\tilde{v}_{3}x_{3}-\tilde{v}_{4}x_{4}).$
Since $\gcd(m_{2},m_{3})=1$, we obtain $m_{2}\mid(a_{1}-1)x_{0}+m_{2}x_{1}$. This implies that $m_{2}\mid(a_{1}-1)x_{0}$. It is not difficult to obtain $\gcd(m_{2},a_{1}-1)=1$ (see equation (3.4.11) in the proof of Theorem 3.1 in [14]), so $m_{2}\mid x_{0}$. On the other hand, since $m_{3}(a_{1}-1)x_{0}\leqslant\tilde{d}=m_{3}m_{2}a_{1}$, we have
$x_{0}\leqslant\dfrac{a_{1}m_{2}}{a_{1}-1}<2m_{2}.$
Since $m_{2}\mid x_{0}$, then $x_{0}=0$ or $x_{0}=m_{2}$.
  • If $x_{0}=m_{2}$, then the equation (3.3.8) is equivalent to
    $m_{3}(a_{1}-1)m_{2}+m_{3}m_{2}x_{1}+m_{2}a_{1}\tilde{v}_{2}x_{2}+m_{2}a_{1}\tilde{v}_{3}x_{3}+m_{2}a_{1}\tilde{v}_{4}x_{4}=m_{3}m_{2}a_{1}$
    Simplifying, we obtain
    $a_{1}(\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4})=m_{3}(1-x_{1}).$
    Since $a_{1}(\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4})\geqslant 0$, we have that $x_{1}=1$ or $x_{1}=0$. If $x_{1}=1$, then $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=0$, which implies that $x_{2}=x_{3}=x_{4}=0$. Thus, we obtain a solution of (3.3.8), which is given by $(m_{2},1,0,0,0)$. On the other hand, if $x_{1}=0$, then $a_{1}(\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4})=m_{3}$, but $\gcd(a_{1},m_{3})=1$ and $a_{1}\neq 1$. Thus, in this case there is no solution.
  • If $x_{0}=0$, then the equation (3.3.8) is given by
    $m_{3}m_{2}x_{1}+m_{2}a_{1}\tilde{v}_{2}x_{2}+m_{2}a_{1}\tilde{v}_{3}x_{3}+m_{2}a_{1}\tilde{v}_{4}x_{4}=m_{3}m_{2}a_{1}.$
    Simplifying, we obtain
    $m_{3}x_{1}=a_{1}(m_{3}-\tilde{v}_{2}x_{2}-\tilde{v}_{3}x_{3}-\tilde{v}_{4}x_{4}).$
    Since $\gcd(a_{1},m_{3})=1$, then we have $m_{3}\mid(m_{3}-\tilde{v}_{2}x_{2}-\tilde{v}_{3}x_{3}-\tilde{v}_{4}x_{4})$. That is, the expression $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}$ can assume two values: $m_{3}$ or 00. First, we suppose that $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}$. This forces $x_{1}=0$. Moreover, since $\tilde{v}_{2}$, $\tilde{v}_{3}$ and $\tilde{v}_{4}$ are given in (3.3.4) and these describe the weight vector $\tilde{\bf w}$ in (3.3.5), then the equation $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}$ has exactly three solutions: $(a_{2},1,0),(0,a_{3},1)$ and $(1,0,a_{4})$. Thus, we obtain three solutions for the equation (3.3.8): $(0,0,a_{2},1,0),(0,0,0,a_{3},1)$ and $(0,0,1,0,a_{4})$. Finally, if we suppose that $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=0$, then $x_{1}=a_{1}$. In this case the only solution is $(0,a_{1},0,0,0)$.
Summarizing, when $a_{1}>2$, we obtain five solutions for Equation (3.3.8):
$(m_{2},1,0,0,0),(0,0,a_{2},1,0),(0,0,0,a_{3},1),(0,0,1,0,a_{4})\mbox{ and }(0,a_{1},0,0,0).$
Proposition 3.3.1.
Let $f$ be and invertible polynomial as in (3.3.1) with weights satisfying conditions (3.2.1). If its Berglund-Hübsch transpose $f^{T}$ with associated weight vector $\bf\tilde{w}$ satisfies the conditions given in the Lemma 3.3.1. Then the complex dimension of the local moduli of the orbifold $Z_{f^{T}}=\{f^{T}=0\}/\mathbb{C}^{*}(\bf{\tilde{w}})$ is zero. Moreover, each $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ is generated by $z_{i}$ for $i=0,\ldots,4.$
Proof.
This result follows from Lemma 3.3.1 and the fact that $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\geqslant 1$ for each $i$ and $\sum_{i=1}^{4}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\leqslant 5.$
Lemma 3.3.2.
Let $f^{T}$ be an invertible polynomial described in (3.3.3) with associated weight vector $\tilde{\bf w}$ as in (3.3.5). If we add the additional condition $a_{1}=2$, then Equation (3.3.7) has six solutions. Moreover, the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(\tilde{d}))$ is given by the set
$\{z_{0}^{m_{2}}z_{1},z_{0}^{2m_{2}},z_{1}^{2},z_{3}z_{2}^{a_{2}},z_{4}{z_{3}}^{a_{3}},z_{2}z_{4}^{a_{4}}\}.$
Proof.
When $a_{1}=2$, Equation (3.3.7) gives
$m_{3}x_{0}+m_{3}m_{2}x_{1}+2m_{2}\tilde{v}_{2}x_{2}+2m_{2}\tilde{v}_{3}x_{3}+2m_{2}\tilde{v}_{4}x_{4}=2m_{3}m_{2},$ (3.3.9)
which can be written as
$m_{3}(x_{0}+m_{2}x_{1})=2m_{2}(m_{3}-\tilde{v}_{2}x_{2}-\tilde{v}_{3}x_{3}-\tilde{v}_{4}x_{4}).$
As $\gcd(m_{2},m_{3})=1$, then $m_{2}\mid(x_{0}+m_{2}x_{1})$. This implies that $m_{2}\mid x_{0}$. Moreover, from (3.3.9) we have $m_{3}x_{0}\leqslant 2m_{3}m_{2}$. Since $m_{3}\neq 0$, then $x_{0}\leqslant 2m_{2}$. Thus, $x_{0}$ can take three values: 00, $m_{2}$ or $2m_{2}$.
Now, we consider two situations: $m_{3}$ is odd or $m_{3}$ is even.
  • a)
    When $m_{3}$ is odd. As $\gcd(a_{1},m_{3})=1$, then the method to find the solutions is similar to what we have done for $x_{0}=0$ or $x_{0}=m_{2}$ in the Lemma 3.3.1. For the additional case $x_{0}=2m_{2}$, we have
    $2m_{3}m_{2}+m_{3}m_{2}x_{1}+2m_{2}\tilde{v}_{2}x_{2}+2m_{2}\tilde{v}_{3}x_{3}+2m_{2}\tilde{v}_{4}x_{4}=2m_{3}m_{2}$
    which has solution $x_{1}=x_{2}=x_{3}=x_{4}=0$. So in this case, we add the solution: $(2m_{2},0,0,0,0)$. Thus, we have exactly six solutions.
  • b)
    When $m_{3}$ is even. Here, we can write $m_{3}=2m_{3}^{\prime}$. Replacing this in Equation (3.3.9) and simplifying, we obtain
    $m_{3}^{\prime}x_{0}+m_{3}^{\prime}m_{2}x_{1}+m_{2}\tilde{v}_{2}x_{2}+m_{2}\tilde{v}_{3}x_{3}+m_{2}\tilde{v}_{4}x_{4}=2m_{3}^{\prime}m_{2}.$ (3.3.10)
    This equation can be written as
    $m_{3}^{\prime}(x_{0}+m_{2}x_{1})=m_{2}(2m_{3}^{\prime}-\tilde{v}_{2}x_{2}-\tilde{v}_{3}x_{3}-\tilde{v}_{4}x_{4}).$
    By a similar argument used above, we obtain $m_{2}\mid x_{0}$ and $x_{0}\leqslant 2m_{2}$. This implies that $x_{0}=0$, $x_{0}=m_{2}$ or $x_{0}=2m_{2}$.
    • If $x_{0}=0$, then Equation (3.3.10) can be written as
      $m_{3}^{\prime}m_{2}x_{1}+m_{2}\tilde{v}_{2}x_{2}+m_{2}\tilde{v}_{3}x_{3}+m_{2}\tilde{v}_{4}x_{4}=2m_{3}^{\prime}m_{2}.$ (3.3.11)
      Simplifying the expression above, we arrive at the equality
      $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}^{\prime}(2-x_{1}).$
      We notice that $x_{1}$ can assume three values: $x_{1}=0$, $x_{1}=1$ or $x_{1}=2$.
        When $x_{1}=0$, the equation above is
        $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=2m_{3}^{\prime}=m_{3}.$
        This equation has exactly three solutions: $(a_{2},1,0),(0,a_{3},1)$ and $(1,0,a_{4})$. Thus, the solutions of Equation (3.3.9) are $(0,0,a_{2},1,0),(0,0,0,a_{3},1)$ and $(0,0,1,0,a_{4})$.
        When $x_{1}=1$, Equation (3.3.11) can be simplified:
        $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}^{\prime}$ (3.3.12)
        If there exists a solution $(\hat{x}_{2},\hat{x}_{3},\hat{x}_{4})$ of this Diophantine equation, then $(2\hat{x}_{2},2\hat{x}_{3},2\hat{x}_{4})$, with $\hat{x}_{i}\in\mathbb{Z}_{0}^{+}$, is solution of the equation
        $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=2m_{3}^{\prime}=m_{3}.$
        Nevertheless, the solutions of this last equation are $(a_{2},1,0),(0,a_{3},1)$ and $(1,0,a_{4})$, which implies that $12\hat{x}_{i}=$ 1 for any $i$. That is, $\hat{x}_{i}$ is not integer, which is not possible.
        When $x_{1}=2$, the equation is
        $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=0,$
        which has trivial solution. Therefore, the solutions of Equation (3.3.9) is $(0,2,0,0,0)$.
    • If $x_{0}=m_{2}$, then Equation (3.3.10) can be reduced to
      $m_{3}^{\prime}+m_{3}^{\prime}x_{1}+\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=2m_{3}^{\prime}$ (3.3.13)
      The above equation can be written as
      $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}^{\prime}(1-x_{1}).$
      Then $x_{1}$ can take two values: $x_{1}=0$ or $x_{1}=1$.
        If $x_{1}=0$, then we obtain Equation (3.3.12), which has no solution.
        If $x_{1}=1$, we obtain the equation
        $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=0,$
        which has trivial solution. In this case the solution obtained is $(m_{2},1,0,0,0)$.
    • If $x_{0}=2m_{2}$. Here, Equation (3.3.10) can be reduced to
      $m_{3}^{\prime}m_{2}x_{1}+m_{2}\tilde{v}_{2}x_{2}+m_{2}\tilde{v}_{3}x_{3}+m_{2}\tilde{v}_{4}x_{4}=0$
      which has trivial solution. In this case, the solution of Equation (3.3.9) is given by $(2m_{2},0,0,0,0)$.
    In view of the foregoing discussion, we find six solutions for Equation (3.3.9):
    $(0,0,a_{2},1,0),(0,0,0,a_{3},1),(0,0,1,0,a_{4}),(0,2,0,0,0),(m_{2},1,0,0,0),\mbox{ and }(2m_{2},0,0,0,0).$
Proposition 3.3.2.
Let $f$ be and invertible polynomial as in (3.3.1) with weights satisfying conditions (3.2.1). If its Berglund-Hübsch transpose $f^{T}$ with associated weight vector $\bf\tilde{w}$ satisfies the conditions given in the Lemma 3.3.2. Then the complex dimension of the local moduli of the orbifold $Z_{f^{T}}=\{f^{T}=0\}/\mathbb{C}^{*}(\bf{\tilde{w}})$ is zero. Moreover, $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))$ is generated by $z_{0}^{m_{2}}$ and $z_{1}$ while for $i\not=1$ $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ is generated by $z_{i}.$
Proof.
First, we will prove that $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))\geqslant 2$. In fact, it is equivalent to show that the quantity of solutions of the following Diophantine equation is not less that two:
$\tilde{w}_{0}x_{0}+\tilde{w}_{1}x_{1}+\tilde{w}_{2}x_{2}+\tilde{w}_{3}x_{3}+\tilde{w}_{4}x_{4}=\tilde{w}_{1}.$
Using the expression for $\bf\tilde{w}$ given in (3.3.5) and replacing $a_{1}=2$ in the equation above, we arrive to the equation
$m_{3}x_{0}+m_{3}m_{2}x_{1}+2m_{2}\tilde{v}_{2}x_{2}+2m_{2}\tilde{v}_{2}x_{3}+2m_{2}\tilde{v}_{4}x_{4}=m_{3}m_{2}.$
Here, we can exhibit at least two solutions: $(m_{2},0,0,0,0)$ and $(0,1,0,0,0)$. Therefore, we obtain $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))\geqslant 2$.
On the other hand, we know that $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\geqslant 1$ for $i\neq 2$. This implies that
$66\leqslant\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\leqslant\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))=$ 6 .
From here, we can conclude that $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))$ has dimension 2 and the set of generators is $\{z_{0}^{m_{2}},z_{1}\}$ and each $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ is one dimensional and it is generated by $z_{i}$ for $i\not=1.$
$\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))-\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))=0.$
Next we discuss the moduli of the orbifold associated to the polynomials considered in this subsection.
Remark 3.3.1.
For $\mathbb{P}(\mathbf{\tilde{w}})=\operatorname{Proj}(S(\mathbf{\tilde{w}}))$, the complex automorphisms of the graded ring $S(\mathbf{\tilde{w}})$ with can be defined on generators for the two types of polynomials given in this subsection thanks to Propositions 3.3.1 and 3.3.2. Notice that the generating set of $H^{0}(\mathbb{P}(\mathbf{\tilde{w}}),\mathcal{O}({\tilde{d}}))$ in all these cases is given by quasi-smooth monomials as in the previous cases, so we can describe the local moduli of the corresponding orbifold, a $\mathbb{Q}$-homology projective 3-space, as done previously. Indeed, let $\alpha_{i},\beta_{1}\in\mathbb{C}^{*},$ then we have
  • If $f$ is a polynomial as in Lemma 3.3.1, then $G(\mathbf{\tilde{w}})$ is given on generators by
    $\varphi_{\mathrm{w}}\left(\begin{array}[]{l}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\end{array}\right)=\left(\begin{array}[]{c}\alpha_{0}z_{0}\\ \alpha_{1}z_{1}\\ \alpha_{2}z_{2}\\ \alpha_{3}z_{3}\\ \alpha_{4}z_{4}\end{array}\right)$
    and the moduli of the orbifold $Z_{f^{T}}$ can be expressed by the quotient
    $Span\{z_{0}^{m_{2}}z_{1},z_{1}^{a_{1}},z_{3}z_{2}^{a_{2}},{z_{4}}z_{3}^{a_{3}},z_{2}z_{4}^{a_{4}}\}/G(\mathbf{\tilde{w}})$
    a zero-dimensional quotient space.
  • If $f$ is a polynomial as in Lemma 3.3.2, then $G(\mathbf{\tilde{w}}))$ is given on generators by
    $\varphi_{\mathrm{w}}\left(\begin{array}[]{l}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\end{array}\right)=\left(\begin{array}[]{c}\alpha_{0}z_{0}\\ \alpha_{1}z_{1}+\beta_{1}z_{0}^{m_{2}}\\ \alpha_{2}z_{2}\\ \alpha_{3}z_{3}\\ \alpha_{4}z_{4}\end{array}\right)$
    and the moduli of the orbifold $Z_{f^{T}}$ can be expressed by the quotient
    $Span\left\{z_{0}^{m_{2}}z_{1},z_{0}^{2m_{2}},z_{1}^{2},z_{3}z_{2}^{a_{2}},z_{4}z_{3}^{a_{3}},z_{2}z_{4}^{a_{4}}\right\}/G(\mathbf{\tilde{w}})$
    again a zero-dimensional quotient space.
In [14], Theorem 4.1 we established the existence of 75 new $\mathbb{Q}$-homology 7-spheres admitting Sasaki-Einstein metrics with not well-formed quotients, that is, with some group elements having codimension 1 fixed point sets. All these links, which are listed in a table in the Appendix in [14], are given by polynomials of chain-cycle type of type II with weights satisfying the conditions given in (3.2.1). So we have:
Theorem 3.3.1.
Let $f$ be and invertible polynomial as in (3.3.1) with weights satisfying conditions (3.2.1). If its Berglund-Hübsch transpose $f^{T}$ with associated weight vector $\bf\tilde{w}$ satisfies the conditions given in either Lemma 3.3.1 or Lemma 3.3.2 and additionally $f$ belongs to one of the 236 $\mathbb{Q}$-homology 7-spheres admitting Sasaki-Einstein metrics found in [5] and [14], then the dimension of the local moduli space of Sasaki-Einstein metrics of the $\mathbb{Q}$-homology 7-spheres at $L_{f^{T}}$ equals zero. Thus $\mathbb{Q}$-homology 7-spheres given as links coming from polynomials $f^{T}$ as above do not admit inequivalent families of Sasaki-Einstein structures.
Proof.
This result follows from Propositions 3.3.1 and 3.3.2 ∎
Remark 3.3.2.
From the list of 75 new examples of Sasaki-Einstein $\mathbb{Q}$-homology 7-sphere given in the Appendix of [14], there are 7 weight vectors $\bf\tilde{w}$ that do not satisfy the additional conditions of either Lemma 3.3.1 or Lemma 3.3.2, actually all of the elements of this table satisfy $a_{1}>2$ but $\gcd(m_{3},a_{1})\not=1.$ For these, we have the following table where $\mu_{\mathbb{R}}$ denotes the real dimension of the local moduli of Sasaki-Einstein metrics of $\mathbb{Q}$-homology 7-spheres at $L_{f^{T}}.$
$\bf\tilde{w}$ $\tilde{d}$ $m_{3}$ $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))$ $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ $\mu_{\mathbb{R}}$
(177,295,270,370,70) 1180 118 5 5 0
(52,663,867,1581,153) 3315 65 9 6 6
(148,777,987,1911,63) 3885 185 9 6 6
(86,3655,5185,595,1445) 10965 129 6 5 2
(86,3655,4165,2635,425) 10965 129 6 5 2
(438,4161,6175,133,1577) 12483 657 6 5 2
(438,4161,4693,3097,95) 12483 657 6 5 2
Declarations
  • Funding: The first author received financial support from Pontificia Universidad Católica del Perú through project DGI PI0655.
  • Data Availability: Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
  • Conflict of interests: The authors have no competing interests to declare that are relevant to the content of this article.